### v1

# The incipient infinite cluster of the FK-Ising model in dimensions $d\geq 3$ and the susceptibility of the high-dimensional Ising model

Abstract
[…]
We then study the susceptibility $\chi(\beta)$ of the nearest-neighbour Ising model on $\mathbb{Z}^{d}$. When $d>4$, we improve a previous result of Aizenman (Comm. Math. Phys., 1982) to obtain the existence of $A>0$ such that, for $\beta<\beta_{c}$,
$\chi(\beta)=\frac{A}{1-\beta/\beta_{c}}(1+o(1)),$
where $o(1)$ tends to 00 as $\beta$ tends to $\beta_{c}$. Additionally, we relate the constant $A$ to the incipient infinite cluster of the double random current.
1 Introduction
In the context of independent bond percolation on the hypercubic lattice $\mathbb{Z}^{d}$ in dimensions $d\geq 2$, it is believed [^2] The absence of infinite cluster at the critical point was successfully proved when $d=2$ [Kes80], and when $d>10$ [HS90a, Har08, FvdH17]. Proving the corresponding result when $3\leq d\leq 10$ is one of the main open problems in percolation theory. that there is no infinite cluster at the critical threshold $p_{c}$. However, the expected size of the cluster of the origin (or susceptibility) is infinite (see [AN84]). This implies that, from the perspective of an observer at the origin, increasingly large (but finite) clusters emerge as larger length scales are considered.
The incipient infinite cluster (IIC) was introduced in the physics literature [AO82, LS83, SC83] as a way to analyse how this extensive structure “emerges” at the critical point. It can be viewed as a description of the not fully materialised infinite cluster. As explained in [Aiz97], a rigorous definition of this object is not clear, and one may obtain different results depending on the chosen approach.
Various methods have been suggested to construct the IIC. In a seminal paper, Kesten [Kes86a] proposed two ways of constructing the IIC in the case of planar Bernoulli percolation. One consists in conditioning the critical measure $\mathbb{P}_{p_{c}}$ on the one-arm event $\{0\xleftrightarrow{}\Lambda_{n}^{c}\}$, where $\Lambda_{n}=[-n,n]\cap\mathbb{Z}^{d}$, and to send $n$ to infinity. The other one consists in conditioning 00 to lie in an infinite cluster in the supercritical measure $\mathbb{P}_{p}$ with $p>p_{c}$, and to send the parameter $p$ to $p_{c}$. Both approaches were shown to converge (weakly) to the same percolation measure, which additionally satisfies almost surely the property that 00 lies in an infinite cluster. The cluster of the origin under this measure is a natural candidate for the IIC. Chayes, Chayes, and Durett [CCD87] proposed a different construction based on the introduction of a well-chosen non-homogeneous bond parameter $p$. Subsequent works by Járai [Jár03a, Jár03b] provided alternative but equivalent constructions of Kesten’s infinite incipient cluster, thus comforting the robustness of the two dimensional IIC for independent bond percolation.
[…]
Fine properties of the IIC were derived using the aforementioned techniques: sub-diffusivity of the random-walk on the IIC has been studied in [Kes86b, BJKS08, KN09], the connection to invasion percolation has been explored in [Jár03b, HJ04], and the relation of the scaling limit of the high-dimensional IIC to the super-Brownian excursion has been investigated in [HS00a, HS00b].
Removing independency makes the problem more intricate. In the setup of the planar random cluster model (or Fortuin–Kasteleyn (FK) percolation) with cluster weight ${1<q\leq 4}$ ($q=1$ corresponds to Bernoulli percolation), the Russo–Seymour–Welsh (RSW) theory [Rus78, SW78] is sufficient to extend Kesten’s argument. See [Kes86a, Jár03b, BS17]. In higher dimensions, the lace expansion (crucial to build the IIC for Bernoulli percolation) has only been developed for $q\in\{1,2\}$. Setting $q=2$ corresponds to the so-called FK-Ising model. However, the lace expansion of the Ising model is developed at the level of the spin model and very little information is available on the percolation side. In particular, it is not clear how to extend the approach of [HJ04]. Nevertheless, other tools are available for the analysis of the case $q=2$.
The goal of this paper is twofold. First, we study the existence of the IIC measure of the FK-Ising model in dimensions $d\geq 3$. Using the relation of the FK-Ising model to the random current representation of the Ising model established in [LW16] (see also [ADCTW19]), together with the mixing property recently obtained in [ADC21], we construct in two different ways a natural candidate for the IIC measure of the FK-Ising model in dimensions $d\geq 3$. Our definitions are similar to the ones suggested in [HJ04]. Along the way, we also construct the corresponding object for the percolation measure induced by random currents.
Second, we study the high-dimensional susceptibility of the Ising model and obtain its exact asymptotics as $\beta\nearrow\beta_{c}$, i.e. we prove that the limit $\lim_{\beta\nearrow\beta_{c}}\chi(\beta)(1-\beta/\beta_{c})$ exists. This improves a result of Aizenman [Aiz82], where up-to-constants estimates were derived. We then relate this limit to the IIC measure of the double random current.
We stress that our proofs do not rely on the lace expansion.
1.1 Definition of the model
The Ising model is one of the most studied models in statistical mechanics (see [DC22] for a review). It is formally defined as follows: given $\Lambda\subset\mathbb{Z}^{d}$ finite, $\tau\in\{-1,0,+1\}^{\mathbb{Z}^{d}}$, and an inverse temperature $\beta\geq 0$, construct a probability measure on $\{-1,+1\}^{\Lambda}$ according to the formula
$\langle F(\sigma)\rangle_{\Lambda,\beta}^{\tau}:=\frac{1}{\mathbf{Z}_{\Lambda,\beta}^{\tau}}\sum_{\sigma\in\{-1,+1\}^{\Lambda}}F(\sigma)\exp\Big{(}\beta\sum_{\begin{subarray}{c}x,y\in\Lambda\\ x\sim y\end{subarray}}\sigma_{x}\sigma_{y}+\beta\sum_{\begin{subarray}{c}x\in\Lambda\\ y\notin\Lambda\\ x\sim y\end{subarray}}\sigma_{x}\tau_{y}\Big{)},$ (1.1)
where ${F:\{-1,+1\}^{\Lambda}\rightarrow\mathbb{R}}$, $\mathbf{Z}_{\Lambda,\beta}^{\tau}$ is the partition function of the model which guarantees that $\langle 1\rangle_{\Lambda,\beta}^{\tau}=1$, and ${x\sim y}$ means that ${|x-y|_{2}=1}$ (where $|\cdot|_{2}$ is the $\ell^{2}$ norm on $\mathbb{R}^{d}$). This definition corresponds to the nearest-neighbour ferromagnetic Ising model on $\Lambda$ with boundary condition $\tau$. When $\tau\equiv 1$ (resp $\tau\equiv 0$), we write $\langle\cdot\rangle^{\tau}_{\Lambda,\beta}=\langle\cdot\rangle^{+}_{\Lambda,\beta}$ (resp. $\langle\cdot\rangle_{\Lambda,\beta}$).
It is a well-known fact (see [Gri67a, Gri67b]) that the measures $\langle\cdot\rangle_{\Lambda,\beta}^{+}$ and $\langle\cdot\rangle_{\Lambda,\beta}$ admit weak limits when $\Lambda\nearrow\mathbb{Z}^{d}$. We respectively denote them by $\langle\cdot\rangle_{\beta}^{+}$ and $\langle\cdot\rangle_{\beta}$. When $d\geq 2$, the model undergoes a phase transition for the vanishing of the magnetisation at a critical parameter $\beta_{c}\in(0,\infty)$ defined by
$\beta_{c}:=\inf\Big{\{}\beta\geq 0,\>m^{*}(\beta):=\langle\sigma_{0}\rangle_{\beta}^{+}>0\Big{\}}.$ (1.2)
The phase transition is continuous in the sense that $m^{*}(\beta_{c})=0$, see [AF86, Wer09, ADCS15, DC19].
The Ising model is classically related to a dependent percolation model: the FK-Ising model, which we now define. For more information we refer to the monograph [Gri06] or the lecture notes [DC19].
Define $\mathbb{E}:=\{\{x,y\}:x\sim y,\>x,y\in\mathbb{Z}^{d}\}$. Let $G=(V,E)$ be a finite subgraph of $\mathbb{Z}^{d}$. Fix $\xi\in\{0,1\}^{\mathbb{E}\setminus E}$. For a fixed percolation configuration $\omega\in\{0,1\}^{E}$, we let $k^{\xi}(\omega)$ be the number of connected components of the graph with vertex set $\mathbb{Z}^{d}$ and edge set $\omega\vee\xi$ that intersect $G$. The FK-Ising model at inverse temperature $\beta\geq 0$ with boundary condition $\xi$ is the measure on $\{0,1\}^{E}$ defined by
$\phi^{\xi}_{G,\beta}[\omega]:=\frac{2^{k^{\xi}(\omega)}\prod_{\{x,y\}\in E}(e^{2\beta}-1)^{\omega(\{x,y\})}}{\mathbf{Z}_{G,\beta,\xi}^{{\rm FK}}},$ (1.3)
where $\mathbf{Z}_{G,\beta,\xi}^{{\rm FK}}$ is the partition function of the model, which ensures that $\phi^{\xi}_{G,\beta}$ is a probability measure. We let $\phi^{1}_{G,\beta}$ and $\phi_{G,\beta}^{0}$ respectively denote the cases $\xi\equiv 1$ and $\xi\equiv 0$. It is classical that the Ising model and the FK-Ising model can be coupled together. This leads to the following formulas: for $x,y\in G$,
$\langle\sigma_{x}\sigma_{y}\rangle_{G,\beta}=\phi_{G,\beta}^{0}[x\xleftrightarrow{}y],\quad\langle\sigma_{x}\sigma_{y}\rangle_{G,\beta}^{+}=\phi_{G,\beta}^{1}[x\xleftrightarrow{}y],\quad\langle\sigma_{x}\rangle_{G,\beta}^{+}=\phi^{1}_{G,\beta}[0\xleftrightarrow{}\partial G],$ (1.4)
where $\partial G$ is the vertex boundary of $G$ defined by $\partial G:=\{x\in G:\exists y\notin G,\>x\sim y\}$. Once again, it is possible to construct (weak) limits of these measures when $G\nearrow\mathbb{Z}^{d}$. We denote them by $\phi^{1}_{\beta}$ and $\phi^{0}_{\beta}$. In particular, $m^{*}(\beta)=\phi^{1}_{\beta}[0\xleftrightarrow{}\infty]$. The FK-Ising model undergoes a phase transition for the existence of an infinite cluster at the parameter $\beta_{c}$ introduced above. That is, an alternative definition of $\beta_{c}$ is
$\beta_{c}=\inf\Big{\{}\beta\geq 0:\phi^{1}_{\beta}[0\xleftrightarrow{}\infty]>0\Big{\}}.$ (1.5)
It is known (see [Bod06, ADCS15, Rao20]) that $\phi_{\beta}^{1}=\phi_{\beta}^{0}$ for all $\beta>0$. We denote the common measure $\phi_{\beta}$. The relation described above ensures that there is no infinite cluster at criticality: $\phi_{\beta_{c}}[0\xleftrightarrow{}\infty]=0$. As for Bernoulli percolation, defining a notion of critical measure conditioned on the event $\{0\xleftrightarrow{}\infty\}$ is therefore not clear.
1.2 Construction of the incipient infinite cluster
The construction of the IIC for the FK-Ising model when $d=2$ is a classical consequence of RSW theory, see [DCKK+20] and references therein. Our first result provides two equivalent constructions of the IIC measure in dimensions $d\geq 3$. We begin by introducing some useful terminology. Let $\mathcal{F}_{0}$ be the set of cylinder events on $\{0,1\}^{\mathbb{E}}$, i.e. the set of local events that are measurable with respect to the state of finitely many edges. Let $\mathcal{F}$ be the $\sigma$-algebra generated by $\mathcal{F}_{0}$.
Theorem 1.1.
Let $d\geq 3$. For all $E\in\mathcal{F}_{0}$, the limit
[…]
exists independently of the manner in which $x$ goes to infinity. We denote it by $\phi^{\infty}[E]$. Moreover, $\phi^{\infty}$ extends to a probability measure on $\mathcal{F}$, that we still denote $\phi^{\infty}$, and which satisfies
$\phi^{\infty}[0\xleftrightarrow{}\infty]=1.$ (1.7)
Note that the above measure is not translation invariant. The cluster of the origin under the measure $\phi^{\infty}$ is a natural candidate for the IIC of the FK-Ising model in dimensions $d\geq 3$. We now provide a subcritical definition of $\phi^{\infty}$ that is inspired by [HJ04].
For $\beta<\beta_{c}$, we define a probability measure $\mathbb{Q}_{\beta}$ on $\{0,1\}^{\mathbb{E}}$ by setting for $E\in\mathcal{F}$,
[…]
where $\chi(\beta):=\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}=\sum_{x\in\mathbb{Z}^{d}}\phi_{\beta}[0\xleftrightarrow{}x]$ is the susceptibility.
Theorem 1.2.
Let $d\geq 3$. For all $E\in\mathcal{F}_{0}$, the limit
$\lim_{\beta\nearrow\beta_{c}}\mathbb{Q}_{\beta}[E]$ (1.9)
exists. We denote it by $\mathbb{Q}_{\beta_{c}}[E]$. Moreover, one has $\mathbb{Q}_{\beta_{c}}[E]=\phi^{\infty}[E]$.
It is not clear how to construct the IIC with Kesten’s approach. We discuss this issue in Section 1.4.
[…]
1.3 Susceptibility of the high-dimensional Ising model
We recall that the susceptibility is defined for $\beta<\beta_{c}$ by
$\chi(\beta):=\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}.$ (1.10)
It is well known that the nearest-neighbor Ising model introduced above is reflection positive, see [FSS76, FILS78, Bis09]. This property provides two tools that are fundamental in the study of the Ising model: the infrared bound [FSS76, FILS78] and the Messager–Miracle-Solé (MMS) inequalities [MMS77]. These tools combined together imply the existence of $C=C(d)>0$ such that for all $\beta\leq\beta_{c}$, and all $x\in\mathbb{Z}^{d}\setminus\{0\}$,
$\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\leq\frac{C}{|x|^{d-2}},$ (1.11)
see [Pan23] for more details. When $d>4$, a matching lower bound (with a different constant) was recently derived [DCP24].
In his seminal paper, Aizenman [Aiz82] used (1.11) to obtain the existence of $C>0$ such that for all $\beta<\beta_{c}$,
$\frac{(2d\beta_{c})^{-1}}{1-\beta/\beta_{c}}\leq\chi(\beta)\leq\frac{C}{1-\beta/\beta_{c}}.$ (1.12)
In fact, for every $\varepsilon>0$, provided that $\beta$ is sufficiently close to $\beta_{c}$, $C$ can be chosen [^3] In fact, it is possible to extract from their methods a slightly better bound, where the bubble is replaced by an open bubble diagram. See discussion below Theorem 1.4 and Remark 4.4. [AG83] as
$C=\frac{1+(2d\beta_{c})B(\beta_{c})}{2d\beta_{c}}(1+\varepsilon),$ (1.13)
where
$B(\beta):=\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}^{2}$ (1.14)
is the so-called bubble diagram. Note that it is finite when $d>4$ by (1.11). Our next result is a strengthening of (1.12).
Theorem 1.3.
Let $d>4$. There exists $A=A(d)>0$ such that, for $\beta<\beta_{c}$,
$\chi(\beta)=\frac{A}{1-\beta/\beta_{c}}(1+o(1)),$ (1.15)
where $o(1)$ tends to 00 as $\beta$ tends to $\beta_{c}$.
This result is new, even in the setup where the lace expansion can be applied [^4] The existence of $A$ has been derived using the lace expansion for the self-avoiding walk model, see for instance [CLS07]. To the best of our knowledge, the exact asymptotic of the susceptibility is not known for Bernoulli percolation. It would be interesting to see whether it is related the IIC as in the case of the FK-Ising model.. We write the asymptotic in terms of $\beta/\beta_{c}$ rather than $\beta_{c}-\beta$ as it makes appear a constant $A(d)$ which should tend to $1$ as $d$ tends to $\infty$. This can be justified by looking at the Curie–Weiss model (which corresponds to $d=\infty$), for which it is known that $A=1$ (see for instance [FV17, Chapter 2]). Below, we provide a proof of this convergence result.
In fact, it is possible to relate the constant $A$ obtained in Theorem 1.3 to the IIC of a double random current. Let $\mathbf{e}_{1}$ be the unit vector with first coordinate equal to $1$. Let $\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}$ be the IIC measure defined in Theorem 3.1. Under this measure, 00 lies in an infinite cluster made of an infinite path emerging from the origin and a collection of finite loops attached to this path. Using the terminology introduced in Section 2, the sources of $\mathbf{n}_{1}$ can be viewed as “$\partial\mathbf{n}_{1}=\{0,\infty\}$”. A similar observation can be made for $\mathbf{e}_{1}$. Hence, under this measure, two infinite structures emerge from two neighbouring points of $\mathbb{Z}^{d}$. The next result relates $A$ to the probability that these clusters avoid each other. If $(\mathbf{n}_{1},\mathbf{n}_{2})\sim\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}$, we let $\mathbf{C}(0)$ (resp. $\mathbf{C}(\mathbf{e}_{1})$) be the (infinite) cluster of 00 (resp. $\mathbf{e}_{1}$) in the percolation configuration induced by $\mathbf{n}_{1}+\mathbf{n}_{2}$.
Theorem 1.4.
Let $d>4$. Let $A$ be the constant of Theorem 1.3. Then,
$A^{-1}=(2d\beta_{c})\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})=\emptyset].$ (1.16)
We obtain Theorems 1.3 and 1.4 using the probabilistic interpretation of the derivative of the susceptibility obtained in [Aiz82], see (4.4). This result already makes appear the probability of occurrence of the event $\{\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})=\emptyset\}$ under the following measure:
$\mathbf{Q}_{\beta}[\cdot]:=\frac{1}{\chi(\beta)^{2}}\sum_{x,y\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\cdot],$ (1.17)
where $\mathbf{P}^{0x,\mathbf{e}_{1}y}_{\beta}$ samples independent currents $\mathbf{n}_{1}$ and $\mathbf{n}_{2}$ with respective source sets $\partial\mathbf{n}_{1}=\{0,x\}$ and $\partial\mathbf{n}_{2}=\{\mathbf{e}_{1},y\}$. In Section 3, we prove that the measure $\mathbf{Q}_{\beta}$ converges weakly as $\beta\nearrow\beta_{c}$ to the natural candidate for the IIC measure of the double random current. We also prove that convergence holds for local events. However, the event of interest is not local. Our main contribution is a localization of the event $\{\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})=\emptyset\}$ which relies on the assumption $d>4$. See Lemma 4.3 for a precise statement.
Limit of $A(d)$ as $d$ tends to infinity
Together with the results of [AG83], we can prove (see Remark 4.4) that
$\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})]\geq\frac{1}{1+(2d\beta_{c})B^{(\mathbf{e}_{1})}(\beta_{c})},$ (1.18)
where
$B^{(\mathbf{e}_{1})}(\beta):=\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{x}\rangle_{\beta}$ (1.19)
is an open bubble diagram. The lace expansion analysis of the Ising model (see [Sak07, Sak22]) implies that
$\lim_{d\rightarrow\infty}2d\beta_{c}=1,\qquad\lim_{d\rightarrow\infty}B^{(\mathbf{e}_{1})}(\beta_{c})=0.$ (1.20)
Together with (1.18), this gives
$\lim_{d\rightarrow\infty}\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})=\emptyset]=1.$ (1.21)
Plugging (1.20) and (1.21) in Theorem 1.4 yields
$\lim_{d\rightarrow\infty}A(d)=1.$ (1.22)
1.4 Open problems
To conclude this introduction, we list a few questions which naturally arise from the above results.
1.4.1 Other definitions of the IIC
In view of the above, it is natural to ask whether Kesten’s approach can be used to construct the IIC of the FK-Ising model. This leads to the following questions.
Open problem 1.
Let $d\geq 3$. Show that one may define the measure $\phi^{\infty}$ by conditioning on the one-arm event:
[…]
In [HvdHH14] this construction was successfully achieved in the context of high-dimensional Bernoulli percolation using the computation of the one-arm exponent [^5] The one-arm exponent $\rho$ of Bernoulli percolation is defined by $\mathbb{P}_{p_{c}}[0\xleftrightarrow{}\partial\Lambda_{n}]\asymp n^{-1/\rho}$. It is known to be equal to $1/2$ when $d>10$. performed in [KN11] (to be more precise, they only need a lower bound on this exponent). To the best of our knowledge, a lower bound on this exponent, even in the mean-field regime, is not known (see [HHS19] for an upper bound).
Open problem 2.
Let $d\geq 3$. Show that one may define the measure $\phi^{\infty}$ from the supercritical regime:
[…]
The proof of Theorem 1.2 relies on the near-critical version of the mixing property of [ADC21]. Such a near-critical tool is not available on the supercritical side and this seems to be the main obstacle to follow this approach.
1.4.2 Behaviour of the susceptibility when $d=4$
In view of Theorem 1.3, it is natural to ask what is the behaviour of the susceptibility when $d=4$, where mean-field behaviour still holds. Renormalization group analysis and universality hypothesis (see for instance [BBS14] for the corresponding result in the setup of the weakly-coupled $\varphi^{4}$ model) suggest the following behaviour.
Conjecture.
Let $d=4$. There exists $A>0$ such that
$\chi(\beta)=\frac{A|\log(1-\beta/\beta_{c})|^{1/3}}{1-\beta/\beta_{c}}(1+o(1)),$ (1.25)
where $o(1)$ tends to 00 as $\beta$ tends to $\beta_{c}$.
The best result regarding this conjecture was obtained in [AG83] where the authors obtained the existence of $C>0$ such that for all $\beta<\beta_{c}$,
$\frac{(2d\beta_{c})^{-1}}{1-\beta/\beta_{c}}\leq\chi(\beta)\leq\frac{C|\log(1-\beta/\beta_{c})|}{1-\beta/\beta_{c}}.$ (1.26)
In the four-dimensional case, Theorem 1.4 suggests that the divergence of $\chi(\beta)(1-\beta/\beta_{c})$ as $\beta$ tends to $\beta_{c}$ is related to the fact that $\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})=\emptyset]=0$. This hints that $d=4$ is (just like for random walks) the critical dimension for the intersection of two neighbouring random current IICs. As we will see in the proof of Theorem 1.4, when $d>4$ it is possible to relate the susceptibility to $\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}(0)\cap\mathbf{C}(\mathbf{e}_{1})\cap\Lambda_{k}=\emptyset]+o(1)$ where $o(1)$ is uniform in $\beta$ and tends to 00 as $k$ tends to infinity. When $d=4$, this “$o(1)$” is not uniform in $\beta$ anymore. One potential idea to tackle this problem is to choose $k=k(\beta)$ as a function of $\beta$. It seems natural to choose $k(\beta)\geq\xi(\beta)$ where we recall that $\xi(\beta)$ is the so-called correlation length, defined by
$\xi(\beta)^{-1}:=\lim_{n\rightarrow\infty}-\frac{1}{n}\log\langle\sigma_{0}\sigma_{n\mathbf{e}_{1}}\rangle_{\beta}.$ (1.27)
Open problem 3.
When $d=4$, relate the asymptotic behaviour of the susceptibility to the probability of avoidance of two neighbouring IICs up to distance $\xi(\beta)^{C}$, where $C\geq 1$.
2 The random current representation
We begin by introducing the random current representation of the Ising model. For more information we refer to the lecture notes [DC19].
2.1 Definition and main properties
Let $\Lambda$ be a finite subset of $\mathbb{Z}^{d}$.
Definition 2.1.
A current $\mathbf{n}$ on $\Lambda$ is a function defined on the edge-set $E(\Lambda):=\{\{x,y\}:\>x,y\in\Lambda\textup{ and }x\sim y\}$ and taking its values in =$\mathbb{N}=\{0,1,\ldots\}$. We let $\Omega_{\Lambda}$ be the set of currents on $\Lambda$. The set of sources of $\mathbf{n}$, denoted by $\partial\mathbf{n}$, is defined as
$\partial\mathbf{n}:=\Big{\{}x\in\Lambda\>:\>\sum_{y\sim x}\mathbf{n}_{x,y}\textup{ is odd}\Big{\}}.$ (2.1)
We also define
$w_{\beta}(\mathbf{n}):=\prod_{\begin{subarray}{c}\{x,y\}\in E(\Lambda)\end{subarray}}\dfrac{\beta^{\mathbf{n}_{x,y}}}{\mathbf{n}_{x,y}!}.$ (2.2)
It is classical [Aiz82, DC19] that the correlation functions of the Ising model are related to currents: for $S\subset\Lambda$, if $\sigma_{S}:=\prod_{x\in S}\sigma_{x}$,
$\left\langle\sigma_{S}\right\rangle_{\Lambda,\beta}=\dfrac{\sum_{\partial\mathbf{n}=S}w_{\beta}(\mathbf{n})}{\sum_{\partial\mathbf{n}=\emptyset}w_{\beta}(\mathbf{n})}.$ (2.3)
The trace of a current $\mathbf{n}$ naturally induces a percolation configuration $(\mathds{1}_{\mathbf{n}_{x,y}>0})_{\{x,y\}\in E(\Lambda)}$ on $\Lambda$. As it turns out, the connectivity properties of this percolation model play a crucial role in the analysis of the Ising model. This motivates the following terminology.
Definition 2.2.
Let $\mathbf{n}\in\Omega_{\Lambda}$ and $x,y\in\Lambda$.
  1. $(i)$
    We say that $x$ is connected to $y$ in $\mathbf{n}$ and write $x\overset{\mathbf{n}}{\longleftrightarrow}y$, if there is a sequence of points $x_{0}=x,x_{1},\ldots,x_{m}=y$ such that $\mathbf{n}_{x_{i},x_{i+1}}>0$ for $10\leq i\leq m-$ 1.
  2. $(ii)$
    The cluster of $x$, denoted by $\mathbf{C}_{\mathbf{n}}(x)$, is the set of points connected to $x$ in $\mathbf{n}$.
The main interest of the random current representation lies in the following combinatorial tool, called the switching lemma, which was first introduced in [GHS70]. The probabilistic picture attached to it was developed in [Aiz82].
Lemma 2.3 (Switching Lemma).
For every $S_{1},S_{2}\subset\Lambda$ and every function $F$ from $\Omega_{\Lambda}$ into $\mathbb{R}$,
$\sum_{\begin{subarray}{c}\mathbf{n}_{1}\in\Omega_{\Lambda}:\>\partial\mathbf{n}_{1}=S_{1}\\ \mathbf{n}_{2}\in\Omega_{\Lambda}:\>\partial\mathbf{n}_{2}=S_{2}\end{subarray}}F(\mathbf{n}_{1}+\mathbf{n}_{2})w_{\beta}(\mathbf{n}_{1})w_{\beta}(\mathbf{n}_{2})\\ =\sum_{\begin{subarray}{c}\mathbf{n}_{1}\in\Omega_{\Lambda}:\>\partial\mathbf{n}_{1}=S_{1}\Delta S_{2}\\ \mathbf{n}_{2}\in\Omega_{\Lambda}:\>\partial\mathbf{n}_{2}=\emptyset\end{subarray}}F(\mathbf{n}_{1}+\mathbf{n}_{2})w_{\beta}(\mathbf{n}_{1})w_{\beta}(\mathbf{n}_{2})\mathds{1}_{(\mathbf{n}_{1}+\mathbf{n}_{2})\in\mathcal{F}_{S_{2}}},$ (2.4)
where $S_{1}\Delta S_{2}=(S_{1}\cup S_{2})\setminus(S_{1}\cap S_{2})$ is the symmetric difference of sets and $\mathcal{F}_{S_{2}}$ is given by
$\mathcal{F}_{S_{2}}=\{\mathbf{n}\in\Omega_{\Lambda}\>,\>\exists\mathbf{m}\leq\mathbf{n}\>,\>\partial\mathbf{m}=S_{2}\}.$ (2.5)
If $S\subset\Lambda$, define a probability measure $\mathbf{P}_{\Lambda,\beta}^{S}$ on $\Omega_{\Lambda}$ as follows: for every $\mathbf{n}\in\Omega_{\Lambda}$,
$\mathbf{P}_{\Lambda,\beta}^{S}[\mathbf{n}]:=\mathds{1}_{\partial\mathbf{n}=S}\frac{w_{\beta}(\mathbf{n})}{Z_{\Lambda,\beta}^{S}},$ (2.6)
where $Z_{\Lambda,\beta}^{S}:=\sum_{\partial\mathbf{n}=S}w_{\beta}(\mathbf{n})$ is a normalisation constant. Moreover, for $S_{1},\ldots,S_{k}\subset\Lambda$, define
$\mathbf{P}_{\Lambda,\beta}^{S_{1},\ldots,S_{k}}:=\mathbf{P}_{\Lambda,\beta}^{S_{1}}\otimes\ldots\otimes\mathbf{P}_{\Lambda,\beta}^{S_{k}}.$ (2.7)
When $S=\{x,y\}$, we will write $\mathbf{P}^{xy}_{\Lambda,\beta}$ instead of $\mathbf{P}^{\{x,y\}}_{\Lambda,\beta}$. If $\mathcal{E}\subset\Omega_{\Lambda}$, we will also write $Z_{\Lambda,\beta}^{S}[\mathcal{E}]:=\sum_{\partial\mathbf{n}=S}w_{\beta}(\mathbf{n})\mathds{1}_{\mathbf{n}\in\mathcal{E}}$. As proved in [ADCS15], if $S$ is a finite even (i.e. $|S|$ is even) subset of $\mathbb{Z}^{d}$, the sequence of probability measures $(\mathbf{P}_{\Lambda,\beta}^{S})_{\Lambda\subset\mathbb{Z}^{d}}$ admits a weak limit as $\Lambda\nearrow\mathbb{Z}^{d}$ that we denote by $\mathbf{P}_{\beta}^{S}$.
The following mixing statement will be the main tool to construct the incipient infinite cluster.
Theorem 2.4 (Mixing property of currents).
Let $d\geq 3$ and $s\geq 1$. Let $\varepsilon>0$. For every $n\geq 1$, there exist $N\geq n$ and $\beta(\varepsilon)<\beta_{c}$ such that, for every $1\leq t\leq s$, every $\beta\in[\beta(\varepsilon),\beta_{c}]$, every $x_{i}\in\Lambda_{n}$ and $y_{i}\notin\Lambda_{N}$ $(i\leq t)$, and every events $E$ and $F$ depending on the restriction of $(\mathbf{n}_{1},\ldots,\mathbf{n}_{s})$ to edges with endpoints within $\Lambda_{n}$ and outside $\Lambda_{N}$ respectively,
$\left|\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[E\cap F]-\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[E]\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[F]\right|\leq\varepsilon.$ (2.8)
Furthermore, for every $x_{1}^{\prime},\ldots,x_{t}^{\prime}\in\Lambda_{n}$ and $y_{1}^{\prime},\ldots,y^{\prime}_{t}\notin\Lambda_{N}$, we have that
$\left|\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[E]-\mathbf{P}_{\beta}^{x_{1}y^{\prime}_{1},\ldots,x_{t}y^{\prime}_{t},\emptyset,\ldots,\emptyset}[E]\right|\leq\varepsilon,$ (2.9)
$\left|\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[F]-\mathbf{P}_{\beta}^{x^{\prime}_{1}y_{1},\ldots,x^{\prime}_{t}y_{t},\emptyset,\ldots,\emptyset}[F]\right|\leq\varepsilon.$ (2.10)
We will only use (2.9). However, for sake of completeness, we state and prove the stronger result.
Theorem 2.4 was obtained by Aizenman and Duminil-Copin [ADC21] to derive the marginal triviality of the scaling limits of Ising and $\varphi^{4}$ systems in dimension four. Their result is even stronger in the sense that it is quantitative. However, their argument only works in dimensions $d\geq 4$. This is mainly due to the fact that the infrared bound (1.11) is not sharp when $d=3$. As they mention (see discussion below [ADC21, Theorem 6.4]), it is possible to extend their strategy to the three-dimensional case to the cost of losing the quantitative statement. We provide a full proof in Section 5, which heavily relies on the fact that $m^{*}(\beta_{c})=0$.
2.2 Coupling with the FK-Ising model
The following result can be found in [ADCTW19, Theorem 3.2] and originates from the observation of [LW16].
Proposition 2.5.
Let $G=(V,E)$ be a subgraph of $\mathbb{Z}^{d}$ and let $S\subset G$ be a finite even subset of $G$. Let $\beta>0$. Let $\mathbf{n}$ is distributed according to $\mathbf{P}_{G,\beta}^{S}$. Let $(\omega_{e})_{e\in\mathbb{E}}$ be an independent Bernoulli percolation with parameter $1-\exp(-\beta)$. For each $e\in E$, we define
$\eta_{e}:=\max(\mathds{1}_{\mathbf{n}_{e}>0},\omega_{e}).$ (2.11)
Then, the law of $\eta$ is exactly $\phi_{G,\beta}^{0}[\>\cdot\>|\>\mathcal{F}_{B}]$, where $\mathcal{F}_{S}$ is the event introduced in (2.5). In particular, if $\mathcal{A}$ is an increasing event,
$\mathbf{P}^{S}_{G,\beta}[\mathcal{A}]\leq\phi_{G,\beta}^{0}[\mathcal{A}\>|\>\mathcal{F}_{S}].$ (2.12)
We will apply the above statement to the special cases $S=\emptyset$ or $S=\{x,y\}$ for $x,y\in G$. In the former case the law of $\eta$ is $\phi^{0}_{G,\beta}$, while in the latter case it is $\phi^{0}_{G,\beta}[\>\cdot\>|\>x\xleftrightarrow{}y]$, i.e. the FK-Ising measure conditioned on the event that $x$ is connected to $y$. Finally, when $G=\mathbb{Z}^{d}$, by the results recalled above, $\phi^{0}_{G,\beta}=\phi_{\beta}$.
3 The incipient infinite cluster
In this section, we construct the IIC measure for both random currents and the FK-Ising model.
3.1 Construction of the IIC measure for random currents
We will prove the following result.
Theorem 3.1 (The IIC measure of the random current).
Let $d\geq 3$ and $s\geq 1$. Let $1\leq t\leq s$. For all $x_{1},\ldots,x_{t}\in\mathbb{Z}^{d}$ there exists a measure $\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}$ (where $\emptyset$ appears $s-t$ times) on $\Omega_{\mathbb{Z}^{d}}$ such that, for all local events $\mathcal{A}\subset(\Omega_{\mathbb{Z}^{d}})^{s}$,
$\lim_{|y_{1}|,\ldots,|y_{t}|\rightarrow\infty}\mathbf{P}_{\beta_{c}}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]=\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[\mathcal{A}],$ (3.1)
regardless of the manner in which $y_{1},\ldots,y_{t}$ are sent to infinity. Moreover, one also has,
$\lim_{\beta\nearrow\beta_{c}}\frac{1}{\chi(\beta)^{t}}\sum_{y_{1},\ldots,y_{t}\in\mathbb{Z}^{d}}\langle\sigma_{x_{1}}\sigma_{y_{1}}\rangle_{\beta}\ldots\langle\sigma_{x_{t}}\sigma_{y_{t}}\rangle_{\beta}\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]=\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[\mathcal{A}].$ (3.2)
Finally, the measure $\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}$ satisfies,
$\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[x_{i}\xleftrightarrow{\mathbf{n}_{i}\>}\infty,\>\forall 1\leq i\leq t]=1.$ (3.3)
Proof.
We let $\mathcal{A}$ be a local event. Let $n\geq 1$ such that $\mathcal{A}$ is measurable in terms of the edges with endpoints within $\Lambda_{n}$.
Let $\varepsilon>0$. Using Theorem 2.4 , we find $N$ large enough such that for all $y_{1},y_{1}^{\prime},\ldots,y_{t},y_{t}^{\prime}\notin\Lambda_{N}$,
$\Big{|}\mathbf{P}_{\beta_{c}}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]-\mathbf{P}_{\beta_{c}}^{x_{1}y_{1}^{\prime},\ldots,x_{t}y_{t}^{\prime},\emptyset,\ldots,\emptyset}[\mathcal{A}]\Big{|}\leq\varepsilon.$ (3.4)
This shows that the sequence $(\mathbf{P}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}_{\beta_{c}}[\mathcal{A}])_{y_{1},\ldots,y_{t}}$ is Cauchy [^6] By this, we mean that for any enumerations $r^{(i)}$, ${1\leq i\leq t}$, of $\mathbb{Z}^{d}$, the sequence $(\mathbf{P}^{x_{1}r^{(1)}_{n},\ldots,x_{t}r^{(t)}_{n},\emptyset,\ldots,\emptyset}_{\beta_{c}}[\mathcal{A}])_{n\geq 1}$ is Cauchy, and all the limits are the same. and thus admits a limit that we denote by $\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[\mathcal{A}]$. This allows to define the measure $\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}$ on the cylinder $\sigma$-algebra of $\Omega_{\mathbb{Z}^{d}}$.
We turn to the proof of (3.2). Let $\varepsilon>0$. By Theorem 2.4 we have the existence of $N$ large enough and $\beta(\varepsilon)$ close enough to $\beta_{c}$, such that for all $\beta\in[\beta(\varepsilon),\beta_{c}]$, for all $y_{1},y_{1}^{\prime},\ldots,y_{t},y_{t}^{\prime}\notin\Lambda_{N}$,
$\Big{|}\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]-\mathbf{P}_{\beta}^{x_{1}y_{1}^{\prime},\ldots,x_{t}y_{t}^{\prime},\emptyset,\ldots,\emptyset}[\mathcal{A}]\Big{|}\leq\frac{\varepsilon}{3}.$ (3.5)
Using (3.1), we can find $z_{1},\ldots,z_{n}\notin\Lambda_{N}$ such that
$\Big{|}\mathbf{P}_{\beta_{c}}^{x_{1}z_{1},\ldots,x_{t}z_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]-\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[\mathcal{A}]\Big{|}\leq\frac{\varepsilon}{3}.$ (3.6)
Finally, we use [ADCS15, Theorem 2.3] to argue that the map $\beta\mapsto\mathbf{P}_{\beta}^{x_{1}z_{1},\ldots,x_{t}z_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]$ is continuous at $\beta_{c}$. This yields the existence of $\beta_{1}\in[\beta(\varepsilon),\beta_{c})$ such that for all $\beta\in[\beta_{1},\beta_{c}]$,
$\Big{|}\mathbf{P}_{\beta}^{x_{1}z_{1},\ldots,x_{t}z_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]-\mathbf{P}_{\beta_{c}}^{x_{1}z_{1},\ldots,x_{t}z_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]\Big{|}\leq\frac{\varepsilon}{3}.$ (3.7)
Combining the last three displayed equations yields: for all $\beta\in[\beta_{1},\beta_{c}]$, for all $y_{1},\ldots,y_{t}\notin\Lambda_{N}$,
$\Big{|}\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]-\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[\mathcal{A}]\Big{|}\leq\varepsilon.$ (3.8)
In particular,
$\limsup_{\beta\nearrow\beta_{c}}\frac{1}{\chi(\beta)^{t}}\sum_{y_{1},\ldots,y_{t}\in\mathbb{Z}^{d}}\Big{(}\prod_{i=1}^{t}\langle\sigma_{x_{i}}\sigma_{y_{i}}\rangle_{\beta}\Big{)}\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset}[\mathcal{A}]\leq\varepsilon+\mathbf{P}^{x_{1}\infty,\ldots,x_{t}\infty,\emptyset,\ldots,\emptyset}[\mathcal{A}],$ (3.9)
where we used that $\lim_{\beta\nearrow\beta_{c}}\tfrac{\chi_{N}(\beta)}{\chi(\beta)}=0$ with $\chi_{N}(\beta):=\sum_{x\in\Lambda_{N}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}$. Since (3.9) holds for any $\varepsilon>0$ this gives one inequality. The other inequality follows similarly. ∎
The known properties of the spin model shed some light over the IIC of the random current. Below we state a (short) non-exhaustive list of properties one may easily derive in the case of the measure $\mathbf{P}^{0\infty,\emptyset}$.
Proposition 3.2.
Let $d\geq 4$. The measure $\mathbf{P}^{0\infty,\emptyset}$ constructed in Theorem 3.1 satisfies the following properties:
  1. $(i)$
    The cluster $\mathbf{C}(0)$ of 00 under $\mathbf{P}^{0\infty,\emptyset}$ is one ended almost surely.
  2. $(ii)$
    For any $y\in\mathbb{Z}^{d}$, one has
    $\mathbf{P}^{0\infty,\emptyset}[0\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\>}y]=\langle\sigma_{0}\sigma_{y}\rangle_{\beta_{c}}.$ (3.10)
    Moreover, there exist $c,C>0$ such that for every $y\in\mathbb{Z}^{d}$ with $|y|\geq 2$,
    $\frac{C}{|y|^{d-2}}\geq\mathbf{P}^{0\infty,\emptyset}[0\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\>}y]\geq\begin{cases}\displaystyle\frac{c}{|y|^{d-2}}&\text{ if }d\geq 5,\\ \displaystyle\frac{c}{|y|^{2}\log|y|}&\text{ if }d=4.\end{cases}$ (3.11)
We postpone the proof of this result to Section 6. The property $(ii)$ above suggests that, when $d\geq 4$, the IIC of the measure $\mathbf{P}^{0\infty,\emptyset}$ is a two-dimensional object. We can easily prove that the above result also holds when $d=3$ under the assumption [^7] In fact, we even expect $\lim_{|x|\rightarrow\infty}|x|^{1+\varepsilon}\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}=0$ for some $\varepsilon>0$. This is related to the fact that the critical exponent $\eta$ of the two-point function is conjectured to satisfy $\eta>0$ in the $3D$ Ising model. that $\lim_{|x|\rightarrow\infty}|x|\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}=0$.
3.2 Construction of the IIC measure for the FK-Ising model
Theorems 1.1 and 1.2 are simple consequences of Proposition 2.5 and Theorem 3.1.
Proof of Theorem 1.1.
Let $\omega$ be a Bernoulli percolation of parameter $1-\exp(-\beta_{c})$, and write $\mathbb{P}_{1-\exp(-\beta_{c})}$ for its law. Let $E\in\mathcal{F}_{0}$ be measurable in terms of edges in $\Lambda_{n}$. By Proposition 2.5, for all $x\in\mathbb{Z}^{d}$,
[…]
Conditioning on the restriction of $\omega$ to the edges in $\Lambda_{n}$,
[…]
For a fixed configuration $\omega_{n}$, the event $\{(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega_{n}\in E\}$ is local and hence by Theorem 3.1 one has
$\lim_{|x|\rightarrow\infty}\sum_{\omega_{n}\in\{0,1\}^{E(\Lambda_{n})}}$ $\mathbb{P}_{1-\exp(-\beta_{c})}[\omega_{n}]\mathbf{P}_{\beta_{c}}^{0x}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega_{n}\in E]$
$=\sum_{\omega_{n}\in\{0,1\}^{E(\Lambda_{n})}}\mathbb{P}_{1-\exp(-\beta_{c})}[\omega_{n}]\mathbf{P}^{0\infty}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega_{n}\in E]$
$=\Big{(}\mathbf{P}^{0\infty}\otimes\mathbb{P}_{1-\exp(-\beta_{c})}\Big{)}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega\in E].$ (3.14)
[…]
Since this holds for every $E\in\mathcal{F}_{0}$, the proof follows by defining $\phi^{\infty}$ to be the law of a percolation configuration sampled by $\mathbf{P}^{0\infty}$ and sprinkled by an independent Bernoulli percolation of parameter $1-\exp(-\beta_{c})$. ∎
Proof of Theorem 1.2.
Using (1.4) and Proposition 2.5, for every $\beta<\beta_{c}$, for every $E\in\mathcal{F}_{0}$,
[…]
$=\frac{1}{\chi(\beta)}\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\Big{(}\mathbf{P}_{\beta}^{0x}\otimes\mathbb{P}_{1-\exp(-\beta)}\Big{)}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega\in E]$ (3.17)
$=\Big{(}\frac{1}{\chi(\beta)}\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\mathbf{P}_{\beta}^{0x}\Big{)}\otimes\mathbb{P}_{1-\exp(-\beta)}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega\in E]$ (3.18)
Once again, if $E$ is measurable in terms of edges in $\Lambda_{n}$, for every $\omega_{n}\in\{0,1\}^{E(\Lambda_{n})}$, Theorem 3.1 implies the following convergence,
$\lim_{\beta\nearrow\beta_{c}}\frac{1}{\chi(\beta)}\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\mathbf{P}_{\beta}^{0x}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega_{n}\in E]=\mathbf{P}^{0\infty}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega_{n}\in E].$ (3.19)
Thus, as above,
$\lim_{\beta\nearrow\beta_{c}}\mathbb{Q}_{\beta}[E]=\Big{(}\mathbf{P}^{0\infty}\otimes\mathbb{P}_{1-\exp(-\beta_{c})}\Big{)}[(\mathds{1}_{\mathbf{n}_{e}>0})_{e\in\mathbb{E}}\vee\omega\in E]=\phi^{\infty}[E],$ (3.20)
where we used Theorem 1.1 (or more specifically (3.15)) in the last equality. ∎
4 Near-critical behaviour of the susceptibility when $d>4$: proof of Theorem 1.3
We now turn to the study of the high-dimensional susceptibility of the Ising model. We begin by reminding a classical consequence of the switching lemma.
4.1 Geometric interpretation of $\frac{\mathrm{d}\chi^{-1}}{\mathrm{d}\beta}$
Set $J_{u,v}=\mathds{1}_{|u-v|_{2}=1}$ and let $|J|:=\sum_{x\in\mathbb{Z}^{d}}J_{0,x}=2d$. The following computation goes back to [Aiz82]. We start by writing
$-\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}=\frac{1}{2}\frac{1}{\chi(\beta)^{2}}\sum_{y\in\mathbb{Z}^{d}}\sum_{u,v\in\mathbb{Z}^{d}}J_{u,v}\langle\sigma_{0}\sigma_{y};\sigma_{u}\sigma_{v}\rangle_{\beta}.$ (4.1)
Now, notice that $\langle\sigma_{0}\sigma_{y};\sigma_{u}\sigma_{v}\rangle_{\beta}=\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}+\langle\sigma_{0}\sigma_{v}\rangle_{\beta}\langle\sigma_{u}\sigma_{y}\rangle_{\beta}+U_{4}^{\beta}(0,y,u,v)$, where $U_{4}^{\beta}(0,y,u,v)$ is Ursell’s four-point function and is defined by
$U_{4}^{\beta}(0,y,u,v):=\langle\sigma_{0}\sigma_{y}\sigma_{u}\sigma_{v}\rangle_{\beta}-\langle\sigma_{0}\sigma_{y}\rangle_{\beta}\langle\sigma_{u}\sigma_{v}\rangle_{\beta}-\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{y}\sigma_{v}\rangle_{\beta}-\langle\sigma_{0}\sigma_{v}\rangle_{\beta}\langle\sigma_{y}\sigma_{u}\rangle_{\beta}.$ (4.2)
A classical consequence of the switching lemma [Aiz82] is the following representation of $U_{4}^{\beta}$:
$U_{4}^{\beta}(0,y,u,v)=-2\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0u,vy}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(u)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(v)\neq\emptyset].$ (4.3)
Hence, using (4.1), (4.3), and the symmetries of the model,
$-\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}$ $=|J|-\frac{1}{\chi(\beta)^{2}}\sum_{y,u,v\in\mathbb{Z}^{d}}J_{u,v}\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}\mathbf{P}^{0u,vy}_{\beta}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(u)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(v)\neq\emptyset]$
$=\frac{1}{\chi(\beta)^{2}}\sum_{y,u,v\in\mathbb{Z}^{d}}J_{u,v}\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}\mathbf{P}^{0u,vy}_{\beta}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(u)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(v)=\emptyset]$
$=\frac{1}{\chi(\beta)^{2}}\sum_{x,y\in\mathbb{Z}^{d}}\sum_{\begin{subarray}{c}w\in\mathbb{Z}^{d}\\ |w|_{2}=1\end{subarray}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{w}\sigma_{y}\rangle_{\beta}\mathbf{P}^{0x,wy}_{\beta}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(w)=\emptyset]$
$=\frac{2d}{\chi(\beta)^{2}}\sum_{x,y\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}^{0x,\mathbf{e}_{1}y}_{\beta}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset].$ (4.4)
See Figure 1 for an illustration. One of the contributions of this paper is a precise analysis of the probability appearing in (4.4). The main difficulty to perform it lies in the lack of independence of $\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)$ and $\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})$, which is usually helpful for the purpose of using the switching lemma. To lighten the notations below, we will write for an event $\mathcal{E}$,
$\overline{\mathbf{P}}_{\beta}^{(\ell)}[\mathcal{E}]:=\frac{1}{\chi(\beta)^{2}}\sum_{\begin{subarray}{c}|x|\geq\ell\\ |y|\geq\ell\end{subarray}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}^{0x,\mathbf{e}_{1}y}_{\beta}[\mathcal{E}].$ (4.5)
This quantity corresponds to averaging the probability of occurrence of an event $\mathcal{E}$ with respect to the position of the sources.
Figure 1: A configuration $\mathbf{n}$ which satisfies the event appearing in (4.4). The dotted line highlights the fact that 00 and $\mathbf{e}_{1}$ are not connected in the percolation configuration induced by $\mathbf{n}$.
Remark 4.1.
In (4.4), the contribution of $x$ or $y$ in $\Lambda_{\ell}$ is bounded by
$4d\frac{\chi_{\ell}(\beta),}{\chi(\beta)}.$ (4.6)
This can be made arbitrarily small by choosing $\beta$ close enough to $\beta_{c}$. As a result, as we approach $\beta_{c}$, it is enough to study $\overline{\mathbf{P}}_{\beta}^{(\ell)}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset]$. This justifies the notation introduced in (4.5).
Definition 4.2 (Local event of avoidance).
Let $k\geq 1$. We introduce the local event
$\mathcal{A}_{k}:=\Big{\{}\mathbf{n}\in\Omega_{\mathbb{Z}^{d}},\>\mathbf{C}_{\mathbf{n}_{|\Lambda_{k}}}(0)\cap\mathbf{C}_{\mathbf{n}_{|\Lambda_{k}}}(\mathbf{e}_{1})=\emptyset\Big{\}},$ (4.7)
or, in words, the event that the clusters of 00 and $\mathbf{e}_{1}$ in the restriction of $\mathbf{n}$ to $\Lambda_{k}(0)$ do not intersect.
The following lemma formalises the fact that the event appearing in (4.4) is essentially local around 00 and $\mathbf{e}_{1}$. This is the main technical step of the proof.
Lemma 4.3.
There exists $C=C(d),\eta>0$ such that, for all $\beta\leq\beta_{c}$, for all $k\geq 1$ and $\ell\geq 2k$,
$\left|\overline{\mathbf{P}}_{\beta}^{(\ell)}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset]-\overline{\mathbf{P}}_{\beta}^{(\ell)}[\mathcal{A}_{k}]\right|\leq Ck^{-\eta}.$ (4.8)
Before proving Lemma 4.3, we show how to conclude the proof of Theorems 1.3 and 1.4.
Proof of Theorems 1.3 and 1.4.
Using Theorem 3.1,
$\lim_{\beta\nearrow\beta_{c}}\frac{1}{\chi(\beta)^{2}}\sum_{x,y\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\mathcal{A}_{k}]=\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathcal{A}_{k}].$ (4.9)
Let $\varepsilon>0$. Combining (4.4), Lemma 4.3 and (4.9), if $\beta$ is sufficiently close to $\beta_{c}$ and $\ell\geq 2k$,
$-\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\leq 4d\frac{\chi_{\ell}(\beta)}{\chi(\beta)}+2d\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathcal{A}_{k}]+\frac{C}{k^{\eta}}+\varepsilon.$ (4.10)
Using Remark 4.1, we may choose $\beta$ even closer to $\beta_{c}$ to get
$-\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\leq 2d\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathcal{A}_{k}]+\frac{C}{k^{\eta}}+2\varepsilon.$ (4.11)
so that, integrating between $\beta$ and $\beta_{c}$ and taking $\beta$ to $\beta_{c}$ yields,
$\limsup_{\beta\nearrow\beta_{c}}\big{[}\chi(\beta)(1-\beta/\beta_{c})\big{]}^{-1}\leq(2d\beta_{c})\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathcal{A}_{k}]+\frac{C}{k^{\eta}}+2\varepsilon.$ (4.12)
Letting $k\rightarrow\infty$ and $\varepsilon\rightarrow 0$,
$\limsup_{\beta\nearrow\beta_{c}}\big{[}\chi(\beta)(1-\beta/\beta_{c})\big{]}^{-1}\leq(2d\beta_{c})\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset].$ (4.13)
Similarly, if $\beta$ is close enough to $\beta_{c}$, $\overline{\mathbf{P}}_{\beta}^{(\ell)}[\mathcal{A}_{k}]\geq\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathcal{A}_{k}]-\tfrac{\varepsilon}{2d}$ so that
$-\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\geq 2d\cdot\overline{\mathbf{P}}_{\beta}^{(\ell)}[\mathcal{A}_{k}]-\frac{C}{k^{\eta}}\geq 2d\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathcal{A}_{k}]-\frac{C}{k^{\eta}}-\varepsilon.$ (4.14)
We deduce again that
$\liminf_{\beta\nearrow\beta_{c}}\big{[}\chi(\beta)(1-\beta/\beta_{c})\big{]}^{-1}\geq(2d\beta_{c})\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset].$ (4.15)
The proof follows readily ∎
Remark 4.4.
In fact, the above proof also shows that
$\lim_{\beta\nearrow\beta_{c}}\left|\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\right|=2d\cdot\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset].$ (4.16)
Using [AG83, (3.4)], this gives
$\mathbf{P}^{0\infty,\mathbf{e}_{1}\infty}[\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset]\geq\frac{1}{1+(2d\beta_{c})B(\beta_{c})}.$ (4.17)
However, it is possible to obtain a slightly better bound, replacing $B(\beta_{c})$ by $B^{(\mathbf{e}_{1})}(\beta_{c})$ (defined in (1.19)). We briefly explain how for sake of completeness. The lower bound on $\big{|}\tfrac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\big{|}$ is based on an improved bound on $U_{\beta}^{4}$ [AG83, (4.1’)]: for $x_{1},x_{2},x_{3},x_{4}\in\mathbb{Z}^{d}$,
$|U_{4}^{\beta}(x_{1},x_{2},x_{3},x_{4})|\leq\sum_{u,v\in\mathbb{Z}^{d}}J_{u,v}\langle\sigma_{x_{4}}\sigma_{v}\rangle_{\beta}\langle\sigma_{x_{2}}\sigma_{v}\rangle_{\beta}(\beta J_{u,v})\dfrac{\partial\langle\sigma_{x_{1}}\sigma_{x_{3}}\rangle_{\beta}}{\partial(\beta J_{u,v})}\\ +\langle\sigma_{x_{1}}\sigma_{x_{2}}\rangle_{\beta}\langle\sigma_{x_{1}}\sigma_{x_{3}}\rangle_{\beta}\langle\sigma_{x_{1}}\sigma_{x_{4}}\rangle_{\beta}+\langle\sigma_{x_{3}}\sigma_{x_{1}}\rangle_{\beta}\langle\sigma_{x_{3}}\sigma_{x_{2}}\rangle_{\beta}\langle\sigma_{x_{3}}\sigma_{x_{4}}\rangle_{\beta}.$ (4.18)
As in (4.1) and below, and using Lebowitz’ inequality $U_{4}^{\beta}\leq 0$, we write
$\left|\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\right|=|J|-\frac{1}{2}\frac{1}{\chi(\beta)^{2}}\sum_{x,y,z\in\mathbb{Z}^{d}}J_{y,z}|U_{4}^{\beta}(0,y,x,z)|.$ (4.19)
Injecting (4.18) in (4.19),
$\left|\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\right|$ $\geq|J|-\frac{1}{2}\frac{1}{\chi(\beta)^{2}}\sum_{x,y,z,u,v\in\mathbb{Z}^{d}}J_{y,z}\langle\sigma_{z}\sigma_{v}\rangle_{\beta}\langle\sigma_{y}\sigma_{v}\rangle_{\beta}(\beta J_{u,v})\dfrac{\partial\langle\sigma_{0}\sigma_{x}\rangle_{\beta}}{\partial(\beta J_{u,v})}$ (4.20)
$\qquad-\frac{1}{2}\frac{1}{\chi(\beta)^{2}}\sum_{x,y,z\in\mathbb{Z}^{d}}J_{y,z}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{0}\sigma_{y}\rangle_{\beta}\langle\sigma_{0}\sigma_{z}\rangle_{\beta}$ (4.21)
$\qquad-\frac{1}{2}\frac{1}{\chi(\beta)^{2}}\sum_{x,y,z\in\mathbb{Z}^{d}}J_{y,z}\langle\sigma_{x}\sigma_{0}\rangle_{\beta}\langle\sigma_{x}\sigma_{y}\rangle_{\beta}\langle\sigma_{x}\sigma_{z}\rangle_{\beta}$ (4.22)
$=:|J|-(I)-(II)-(III).$ (4.23)
Now, using translation invariance [^8] In [AG83], the authors apply the Cauchy–Schwarz inequality and get $B(\beta)$ instead of $B^{(\mathbf{e}_{1})}(\beta)$.
$(I)$ $=\frac{1}{2}\frac{1}{\chi(\beta)^{2}}\sum_{x,u,v\in\mathbb{Z}^{d}}(\beta J_{u,v})\dfrac{\partial\langle\sigma_{0}\sigma_{x}\rangle_{\beta}}{\partial(\beta J_{u,v})}\Big{(}\sum_{y\in\mathbb{Z}^{d}}\sum_{|w|_{2}=1}\langle\sigma_{y}\sigma_{v}\rangle_{\beta}\langle\sigma_{y+w}\sigma_{v}\rangle_{\beta}\Big{)}$ (4.24)
$=(2d\beta)B^{(\mathbf{e}_{1})}(\beta)\left|\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\right|,$ (4.25)
where we recall that $B^{(\mathbf{e}_{1})}(\beta)=\sum_{x\in\mathbb{Z}^{d}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{x}\rangle_{\beta}$. Moreover, using the Cauchy–Schwarz inequality
$(II)+(III)\leq\frac{|J|B(\beta)}{\chi(\beta)}.$ (4.26)
Plugging (4.25) and (4.26) in (4.23) gives
$\left|\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\right|\geq\frac{2d}{1+(2d\beta_{c})B^{(\mathbf{e}_{1})}(\beta_{c})}\Big{(}1-\frac{B(\beta_{c})}{\chi(\beta)}\Big{)}.$ (4.27)
Since $B(\beta_{c})<\infty$ by (1.11), one gets
$\lim_{\beta\nearrow\beta_{c}}\left|\frac{\mathrm{d}\chi^{-1}(\beta)}{\mathrm{d}\beta}\right|\geq\frac{2d}{1+(2d\beta_{c})B^{(\mathbf{e}_{1})}(\beta_{c})},$ (4.28)
which immediately yields the improvement on (4.17) by (4.16).
4.2 Proof of Lemma 4.3
The rest of the section is devoted to the proof of the localization result of Lemma 4.3. We will use the notion of backbone of a current introduced in [ABF87, Section 4] (see also [ADCTW19, Pan23]). A current $\mathbf{n}$ on a $\Lambda\subset\mathbb{Z}^{d}$ finite with sources $\partial\mathbf{n}=\{x,y\}$ has at least one path from $x$ to $y$ in the percolation configuration induced by $\mathbf{n}$. The backbone of $\mathbf{n}$ is an exploration of one such path. We fix an arbitrary ordering $\prec$ of the edges of $E(\Lambda)$.
Definition 4.5.
Let $\mathbf{n}\in\Omega_{\Lambda}$ with $\partial\mathbf{n}=\{x,y\}$. The backbone of $\mathbf{n}$, which we denote by $\Gamma(\mathbf{n})$, is the unique oriented and edge self-avoiding path from $x$ to $y$, supported on edges with $\mathbf{n}$ odd, which is minimal for $\prec$. The backbone $\Gamma(\mathbf{n})=\{x_{i}x_{i+1}:0\leq i<k\}$ is obtained via the following procedure:
  1. $(i)$
    Set $x_{0}=x$. The first edge $x_{0}x_{1}$ of $\Gamma(\mathbf{n})$ is the earliest of all the edges $e$ emerging from $x_{0}$ for which $\mathbf{n}_{e}$ is odd. All edges $e\ni x_{0}$ satisfying $e\prec x_{0}x_{1}$ and $e\neq x_{0}x_{1}$ are explored and have $\mathbf{n}_{e}$ even.
  2. $(ii)$
    Each edge $x_{i}x_{i+1}$, $i\geq 1$, is the first of all edges $e$ emerging from $x_{i}$ that have not been explored previously, and for which $\mathbf{n}_{e}$ is odd.
  3. $(iii)$
    The exploration stops when it reaches a vertex from which there are no more odd and non-explored edges available. This must happen at $x_{k}=y$.
Let $\overline{\Gamma}(\mathbf{n})$ denote the set of explored edges, i.e. $\Gamma(\mathbf{n})$ together with all explored even edges.
By definition, $\overline{\Gamma}(\mathbf{n})$ is determined by $\Gamma(\mathbf{n})$ and the ordering $\prec$, and similarly $\Gamma(\mathbf{n})$ can be recovered from $\overline{\Gamma}(\mathbf{n})$ by choosing odd edges in $\overline{\Gamma}(\mathbf{n})$ according to $\prec$. The current $\mathbf{n}\setminus\overline{\Gamma}(\mathbf{n})$, defined to be the restriction of $\mathbf{n}$ to the complement of $\overline{\Gamma}(\mathbf{n})$, is sourceless. Moreover, we can write
$\langle\sigma_{x}\sigma_{y}\rangle_{\Lambda,\beta}=\sum_{\gamma:x\rightarrow y}\rho_{\Lambda}(\gamma),$ (4.29)
where for a path $\gamma:x\rightarrow y$,
$\rho_{\Lambda}(\gamma):=\langle\sigma_{x}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\Lambda,\beta}^{xy}[\Gamma(\mathbf{n})=\gamma].$ (4.30)
When $\Lambda=\mathbb{Z}^{d}$, we write $\rho(\gamma)=\rho_{\Lambda}(\gamma)$. We will use the following useful property of backbones, often referred to as the chain rule (see [ABF87]).
Proposition 4.6 (Chain rule for the backbone).
Let $\Lambda\subset\mathbb{Z}^{d}$. Let $x,y,u,v\in\Lambda$. Then,
$\mathbf{P}^{xy}_{\Lambda,\beta}[\Gamma(\mathbf{n})\textup{ passes through }u\textup{ first and then through }v]\leq\frac{\langle\sigma_{x}\sigma_{u}\rangle_{\Lambda,\beta}\langle\sigma_{u}\sigma_{v}\rangle_{\Lambda,\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\Lambda,\beta}}{\langle\sigma_{x}\sigma_{y}\rangle_{\Lambda,\beta}}.$ (4.31)
In particular, if $u=v$,
$\mathbf{P}^{xy}_{\Lambda,\beta}[\Gamma(\mathbf{n})\textup{ passes through }u]\leq\frac{\langle\sigma_{x}\sigma_{u}\rangle_{\Lambda,\beta}\langle\sigma_{u}\sigma_{y}\rangle_{\Lambda,\beta}}{\langle\sigma_{x}\sigma_{y}\rangle_{\Lambda,\beta}}.$ (4.32)
Proof of Lemma 4.3.
Clearly $\{\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})=\emptyset\}\subset\mathcal{A}_{k}$. It then suffices to analyse the event $\mathcal{B}:=\mathcal{A}_{k}\cap\{\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})\neq\emptyset\}$. Note that if the two clusters intersect, there must be a path from $\mathbf{e}_{1}$ to $\Gamma(\mathbf{n}_{1})$ that does not use any edge of $\overline{\Gamma}(\mathbf{n}_{1})$, except maybe for the last edge:
$\{\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(0)\cap\mathbf{C}_{\mathbf{n}_{1}+\mathbf{n}_{2}}(\mathbf{e}_{1})\neq\emptyset\}\subset\bigcup_{\begin{subarray}{c}u\in\mathbb{Z}^{d}\\ v\in\mathbb{Z}^{d}\\ u\sim v\end{subarray}}\{u\in\Gamma(\mathbf{n}_{1})\}\cap\{\mathbf{e}_{1}\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\>}v\}\cap\{(\mathbf{n}_{1}+\mathbf{n}_{2})_{u,v}>0\}.$ (4.33)
Let $\mathcal{B}_{u,v}:=\mathcal{B}\cap\{u\in\Gamma(\mathbf{n}_{1})\}\cap\{\mathbf{e}_{1}\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\>}v\}\cap\{(\mathbf{n}_{1}+\mathbf{n}_{2})_{u,v}>0\}$. Let $m=k^{\delta}$ for some small $\delta>0$ to be fixed.
$\bullet$ We first look at the case $u\notin\Lambda_{m}$, see Figure 2 for an illustration. Exploring $\Gamma(\mathbf{n}_{1})$, and forgetting about $\mathcal{B}\cap\{(\mathbf{n}_{1}+\mathbf{n}_{2})_{u,v}>0\}$, we get [^9] A similar argument is used in the proof of [Pan23, Lemma 6.13].
$\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\mathcal{B}_{u,v}]$ $\leq\sum_{\begin{subarray}{c}\gamma:0\rightarrow x\\ \gamma\ni u\end{subarray}}\rho(\gamma)\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}^{\emptyset,\mathbf{e}_{1}y}_{\overline{\gamma}^{c},\mathbb{Z}^{d},\beta}[\mathbf{e}_{1}\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\gamma}\>}v]$ (4.34)
$=\sum_{\begin{subarray}{c}\gamma:0\rightarrow x\\ \gamma\ni u\end{subarray}}\rho(\gamma)\langle\sigma_{\mathbf{e}_{1}}\sigma_{v}\rangle_{\overline{\gamma}^{c},\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}$ (4.35)
$\leq\langle\sigma_{\mathbf{e}_{1}}\sigma_{v}\rangle_{\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}\sum_{\begin{subarray}{c}\gamma_{1}:0\rightarrow u\\ \gamma_{2}:u\rightarrow x\end{subarray}}\rho(\gamma_{1}\circ\gamma_{2})$ (4.36)
$\leq\langle\sigma_{\mathbf{e}_{1}}\sigma_{v}\rangle_{\beta}\langle\sigma_{v}\sigma_{y}\rangle_{\beta}\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{u}\sigma_{x}\rangle_{\beta},$ (4.37)
where we used the switching lemma on the second line, Griffiths’ inequality on the third line, and Proposition 4.6 on the last line. As a result, using the infrared bound (1.11), there exists $C_{1}=C_{1}(d)>0$ such that
$\frac{1}{\chi(\beta)^{2}}\sum_{|x|,|y|\geq\ell}\sum_{u\notin\Lambda_{m}}\sum_{v\sim u}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\mathcal{B}_{u,v}]\leq\sum_{u\notin\Lambda_{m}}\sum_{v\sim u}\langle\sigma_{0}\sigma_{u}\rangle_{\beta}\langle\sigma_{\mathbf{e_{1}}}\sigma_{u}\rangle_{\beta}\leq\frac{C_{1}}{m^{d-4}}.$ (4.38)
Figure 2: On the left, an illustration of a configuration realizing $\mathcal{B}_{u,v}$ for $u\notin\Lambda_{m}$. The backbone of $\mathbf{n}_{1}$ (resp. $\mathbf{n}_{2}$) is the black (resp. red) bold line. A string of loops connects $\Gamma(\mathbf{n}_{2})$ to $\Gamma(\mathbf{n}_{1})$ in $\mathbf{n}_{1}+\mathbf{n}_{2}$ without using any edges of $\overline{\Gamma}(\mathbf{n}_{1})$. On the right, a diagrammatic representation of the bound obtained in (4.37).
$\bullet$ We now turn to the case $u\in\Lambda_{m}$. We introduce an additional intermediate scale $M=m^{A}\leq k$ with $A>0$ large enough [^10] How large $A$ can be taken will depend of how small $\delta$ is chosen. to be fixed. Notice that under the occurrence of $\mathcal{B}_{u,v}$, any path connecting $\mathbf{e}_{1}$ to $v$ in the complement of $\overline{\Gamma}(\mathbf{n}_{1})$ has to [^11] This is where the event $\{(\mathbf{n}_{1}+\mathbf{n}_{2})_{u,v}>0\}$ becomes useful in the definition of $\mathcal{B}_{u,v}$. exit the box. We claim that if $\mathcal{B}_{u,v}$ occurs then one of the following events occurs (see Figure 3),
  1. -
    $\mathcal{G}_{1}:$ the backbone $\Gamma(\mathbf{n}_{2})$ does two successive crossings of $\textup{Ann}(M,k)$,
  2. -
    $\mathcal{G}_{2}:$ the event $\mathcal{F}_{1}$ does not occur and there exist $s\in\partial\Lambda_{M}$ and $t\in\partial\Lambda_{m}$ such that $s\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus(\overline{\Gamma}(\mathbf{n}_{1})\cup\overline{\Gamma}(\mathbf{n}_{2}))\>}t$.
Indeed, explore a path from $\mathbf{e}_{1}$ to $v$ in $\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})$. As mentioned above, this path must reach $\partial\Lambda_{M}$ a last time at a vertex $s$ before reaching $v$. If $\mathcal{G}_{1}$ does not occur, then $s$ must be connected to $v$ without using any edges of $\overline{\Gamma}(\mathbf{n}_{2})$. If not, there would be a path inside $\Lambda_{M}$ connecting $s$ to both $v$ and a point [^12] We are using the fact that any edge of $\overline{\Gamma}(\mathbf{n}_{2})$ is incident to a vertex of $\Gamma(\mathbf{n}_{2})$. of $\Gamma(\mathbf{n}_{2})$. Since $(\mathbf{n}_{1}+\mathbf{n}_{2})_{u,v}>0$ this would create a connection between 00 and $\mathbf{e}_{1}$ that is entirely included in $\Lambda_{k}$, which is impossible by assumption.
We begin with a bound of $\mathcal{G}_{1}$. A similar bound was already used in [ADC21, (4.21)]. By the chain rule of Proposition 4.6,
$\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\mathcal{G}_{1}]$ $=\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{\mathbf{e}_{1}y}[\mathcal{G}_{1}]$ (4.39)
$\leq\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\sum_{\begin{subarray}{c}a\in\partial\Lambda_{k}\\ b\in\partial\Lambda_{M}\end{subarray}}\langle\sigma_{\mathbf{e}_{1}}\sigma_{a}\rangle_{\beta}\langle\sigma_{a}\sigma_{b}\rangle_{\beta}\langle\sigma_{b}\sigma_{y}\rangle_{\beta}.$ (4.40)
Using once again the infrared bound (1.11), there exists $C_{2}=C_{2}(d)>0$ such that,
$\overline{\mathbf{P}}^{(\ell)}_{\beta}[\mathcal{G}_{1}]\leq\sum_{\begin{subarray}{c}a\in\partial\Lambda_{k}\\ b\in\partial\Lambda_{M}\end{subarray}}\langle\sigma_{\mathbf{e}_{1}}\sigma_{a}\rangle_{\beta}\langle\sigma_{a}\sigma_{b}\rangle_{\beta}\leq C_{2}\frac{k^{d-1}M^{d-1}}{k^{2d-4}}=C_{2}\frac{M^{d-1}}{k^{d-3}}=C_{2}\frac{k^{(d-1)\delta A}}{k^{d-3}}.$ (4.41)
Figure 3: An illustration of the events $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. The backbone of $\mathbf{n}_{1}$ (resp. $\mathbf{n}_{2}$) is the black (resp. red) bold line. On the left, $\Gamma(\mathbf{n}_{2})$ crosses $\text{Ann}(M,k)$ twice and a string of loops creates a connection between the clusters of 00 and $\mathbf{e}_{1}$. On the right, $\Gamma(\mathbf{n}_{2})$ only crosses $\text{Ann}(M,k)$ once. This forces the existence of a crossing of $\textup{Ann}(m,M)$ in $\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\cup\overline{\Gamma}(\mathbf{n}_{2})$.
We turn to the bound of $\mathcal{G}_{2}$. Write,
$\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\mathcal{G}_{2}]$ $\leq\sum_{\begin{subarray}{c}s\in\partial\Lambda_{M}\\ t\in\partial\Lambda_{m}\end{subarray}}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[s\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus(\overline{\Gamma}(\mathbf{n}_{1})\cup\overline{\Gamma}(\mathbf{n}_{2}))\>}t].$ (4.42)
We claim that $\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[C(s,t)]\leq\langle\sigma_{s}\sigma_{t}\rangle_{\beta}^{2}$ where ${C(s,t):=\{s\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus(\overline{\Gamma}(\mathbf{n}_{1})\cup\overline{\Gamma}(\mathbf{n}_{2}))\>}t\}}$. Indeed, going to partition functions [^13] For full disclosure, this step would require working in a finite volume $\Lambda$ and then taking the limit $\Lambda\nearrow\mathbb{Z}^{d}$. We omit this detail here as it does not affect the argument.
$Z_{\beta}^{0x,\mathbf{e}_{1}y}[C(s,t)]=\sum_{\begin{subarray}{c}\gamma_{1}:0\rightarrow x\\ \gamma_{2}:\mathbf{e}_{1}\rightarrow y\end{subarray}}Z_{\beta}^{0x}[\Gamma(\mathbf{n}_{1})=\gamma_{1}]Z_{\beta}^{\mathbf{e}_{1}y}[\Gamma(\mathbf{n}_{2})=\gamma_{2}]\mathbf{P}^{\emptyset,\emptyset}_{\overline{\gamma}_{1}^{c},\overline{\gamma}_{2}^{c},\beta}[C(s,t)].$ (4.43)
Now, write for $\mathbf{n}_{1}\in\Omega_{\overline{\gamma}_{1}^{c}}$ and $\mathbf{n}_{2}\in\Omega_{\overline{\gamma}_{2}^{c}}$
$\mathbf{n}_{1}=\mathbf{k}_{1}+\mathbf{m}_{1},\qquad\mathbf{k}_{1}(e)=\mathbf{n}_{1}(e)\mathds{1}\{e\in\overline{\gamma}_{2}\}\qquad(e\in\mathbb{E}),$ (4.44)
$\mathbf{n}_{2}=\mathbf{k}_{2}+\mathbf{m}_{2},\qquad\mathbf{k}_{2}(e)=\mathbf{n}_{2}(e)\mathds{1}\{e\in\overline{\gamma}_{1}\}\qquad(e\in\mathbb{E}),$ (4.45)
so that $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ are both currents on $(\overline{\gamma}_{1}\cup\overline{\gamma}_{2})^{c}$. Then,
$Z^{\emptyset,\emptyset}_{\overline{\gamma}_{1}^{c},\overline{\gamma}_{2}^{c},\beta}[C(s,t)]$ $=\sum_{\mathbf{k}_{1},\mathbf{k}_{2}}w_{\beta}(\mathbf{k}_{1})w_{\beta}(\mathbf{k}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{k}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{k}_{2}\end{subarray}}w_{\beta}(\mathbf{m}_{1})w_{\beta}(\mathbf{m}_{2})\mathds{1}\{C(s,t)\}$ (4.46)
$=\sum_{\mathbf{k}_{1},\mathbf{k}_{2}}w_{\beta}(\mathbf{k}_{1})w_{\beta}(\mathbf{k}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{k}_{1}\Delta\{s,t\}\\ \partial\mathbf{m}_{2}=\partial\mathbf{k}_{2}\Delta\{s,t\}\end{subarray}}w_{\beta}(\mathbf{m}_{1})w_{\beta}(\mathbf{m}_{2})\mathds{1}\{C(s,t)\}$ (4.47)
$=Z^{st,st}_{\overline{\gamma}_{1}^{c},\overline{\gamma}_{2}^{c},\beta}[C(s,t)]\leq Z^{st,st}_{\overline{\gamma}_{1}^{c},\overline{\gamma}_{2}^{c},\beta},$ (4.48)
where on the second line we used a slightly more general version of the switching lemma, called the switching principle, see [ADCTW19, Lemma 2.1]. As a result,
$\mathbf{P}^{\emptyset,\emptyset}_{\overline{\gamma}_{1}^{c},\overline{\gamma}_{2}^{c},\beta}[C(s,t)]\leq\langle\sigma_{s}\sigma_{t}\rangle_{\overline{\gamma}_{1}^{c},\beta}\langle\sigma_{s}\sigma_{t}\rangle_{\overline{\gamma}_{2}^{c},\beta}.$ (4.49)
Collecting the above work and using Griffiths’ inequality yields
$\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[C(s,t)]\leq\sum_{\begin{subarray}{c}\gamma_{1}:0\rightarrow x\\ \gamma_{2}:\mathbf{e}_{1}\rightarrow y\end{subarray}}\mathbf{P}_{\beta}^{0x}[\Gamma(\mathbf{n}_{1})=\gamma_{1}]\mathbf{P}_{\beta}^{\mathbf{e}_{1}y}[\Gamma(\mathbf{n}_{2})=\gamma_{2}]\langle\sigma_{s}\sigma_{t}\rangle_{\overline{\gamma}_{1}^{c},\beta}\langle\sigma_{s}\sigma_{t}\rangle_{\overline{\gamma}_{2}^{c},\beta}\leq\langle\sigma_{s}\sigma_{t}\rangle_{\beta}^{2}.$ (4.50)
Injecting (4.50) in (4.42) and averaging over $x$ and $y$ gives,
$\overline{\mathbf{P}}^{(\ell)}_{\beta}[\mathcal{G}_{2}]\leq\sum_{\begin{subarray}{c}s\in\partial\Lambda_{M}\\ t\in\partial\Lambda_{m}\end{subarray}}\langle\sigma_{s}\sigma_{t}\rangle_{\beta}^{2}\leq C_{3}\frac{M^{d-1}m^{d-1}}{M^{2d-4}}=C_{3}\frac{k^{\delta(d-1)}}{k^{A\delta(d-3)}}.$ (4.51)
Combining (4.41) and (4.51) gives $C_{4}=C_{4}(d),C_{5}=C_{5}(d)>0$ such that
$\frac{1}{\chi(\beta)^{2}}\sum_{|x|,|y|\geq\ell}\sum_{u\in\Lambda_{m}}\sum_{v\sim u}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}\langle\sigma_{\mathbf{e}_{1}}\sigma_{y}\rangle_{\beta}\mathbf{P}_{\beta}^{0x,\mathbf{e}_{1}y}[\mathcal{B}_{u,v}]\leq C_{4}m^{d}\left(\frac{k^{\delta A}}{k^{d-3}}+\frac{k^{\delta(d-1)}}{k^{A\delta(d-3)}}\right)\leq C_{5}k^{-\eta},$ (4.52)
for some $\eta>0$ sufficiently small, provided $\delta$ is small enough and $A$ is large enough. The proof follows readily. ∎
5 Proof of the mixing property
In this section, we prove Theorem 2.4. We follow the strategy of [ADC21]. The argument only requires a modification in one of the steps (namely [ADC21, Lemma 6.7]), but we provide a full proof for sake of completeness. One essential ingredient to the proof will be the following result.
Theorem 5.1 (Absence of infinite cluster at $\beta_{c}$, [ADCS15]).
Let $d\geq 3$. Then,
$\phi_{\beta_{c}}[0\xleftrightarrow{}\infty]=0.$ (5.1)
We begin by recalling a fundamental result: the existence of regular scales.
Definition 5.2 (Regular scales).
Fix $c,C>0$. An annular region Ann$(n/2,8n)$ is said to be $(c,C)$-regular if the following properties hold:
  1. $(\mathbf{P1})$
    for every $x,y\in\textup{Ann}(n/2,8n)$, $\langle\sigma_{0}\sigma_{y}\rangle_{\beta}\leq C\langle\sigma_{0}\sigma_{x}\rangle_{\beta}$,
  2. $(\mathbf{P2})$
    for every $x,y\in\textup{Ann}(n/2,8n)$, $|\langle\sigma_{0}\sigma_{x}\rangle_{\beta}-\langle\sigma_{0}\sigma_{y}\rangle_{\beta}|\leq\frac{C|x-y|}{|x|}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}$,
  3. $(\mathbf{P3})$
    $\chi_{2n}(\beta)-\chi_{n}(\beta)\geq c\chi_{n}(\beta)$,
  4. $(\mathbf{P4})$
    for every $x\in\Lambda_{n}$ and $y\notin\Lambda_{Cn}$, $\langle\sigma_{0}\sigma_{y}\rangle_{\beta}\leq\frac{1}{2}\langle\sigma_{0}\sigma_{x}\rangle_{\beta}$.
A scale $k$ is said to be regular if $n=2^{k}$ is such that $\textup{Ann}(n/2,8n)$ is $(c,C)$-regular, a vertex $x\in\mathbb{Z}^{d}$ is said to be in a regular scale if it belongs to an annulus $\textup{Ann}(n,2n)$ with $n=2^{k}$ and $k$ a regular scale.
Theorem 5.3 (Existence of regular scales, [ADC21, Theorem 5.12]).
Let $d\geq 3$. Let $\gamma>2$. There exist $\mathbf{c},\mathbf{C}>0$ such that, for every $\beta\leq\beta_{c}$, and every $1\leq n^{\gamma}\leq N\leq\xi(\beta)$, there are at least $\mathbf{c}\log_{2}\left(\frac{N}{n}\right)$ $(\mathbf{c},\mathbf{C})$-regular scales between $n$ and $N$.
We are now equipped to prove Theorem 2.4. Below, $\mathbf{c},\mathbf{C}$ are given by Theorem 5.3.
Let $d\geq 3$ and $\varepsilon>0$. Fix $n\geq 1$, $s\geq 1$, and $t$ satisfying $1\leq t\leq s$. Let $\beta\leq\beta_{c}$ to be taken sufficiently close to $\beta_{c}$, and let $N\geq n$ to be taken large enough. Introduce integers $m,M$ such that $n\leq m\leq M\leq N$.
For $\mathbf{x}=(x_{1},\ldots,x_{t})$ and $\mathbf{y}=(y_{1},\ldots,y_{t})$, we define
$\mathbf{P}^{\mathbf{xy}}_{\beta}:=\mathbf{P}_{\beta}^{x_{1}y_{1},\ldots,x_{t}y_{t},\emptyset,\ldots,\emptyset},\qquad\mathbf{P}_{\beta}^{\mathbf{xy},\emptyset}:=\mathbf{P}_{\beta}^{\mathbf{xy}}\otimes\mathbf{P}^{\emptyset,\ldots,\emptyset}_{\beta},$ (5.2)
where $\mathbf{P}_{\beta}^{\emptyset,\ldots,\emptyset}$ is the law of a sum of $s$ independent sourceless currents that we denote by $(\mathbf{n}_{1}^{\prime},\ldots,\mathbf{n}_{s}^{\prime})$. We also let $\mathbf{E}_{\beta}^{\mathbf{xy}}$ and $\mathbf{E}_{\beta}^{\mathbf{xy},\emptyset}$ be the expectations with respect to these measures. If $p\geq 1$, define for $y\notin\Lambda_{2dp}$,
$\mathbb{A}_{y}(p):=\left\{u\in\text{Ann}(p,2p)\>:\>\forall x\in\Lambda_{p/d},\>\langle\sigma_{x}\sigma_{y}\rangle_{\beta}\leq\left(1+\mathbf{C}\frac{|x-u|}{|y|}\right)\langle\sigma_{u}\sigma_{y}\rangle_{\beta}\right\}.$ (5.3)
Note that if $y$ is in a regular scale, then $\mathbb{A}_{y}(p)=\text{Ann}(p,2p)$. Moreover, the MMS inequalities ensure that this set is not empty, see [ADC21, Remark 6.5].
Let $\mathfrak{K}$ be the set of regular scales $k$ between $m$ and $M/2$ with the property that for all $k<k^{\prime}\in\mathfrak{K}$, $2^{k^{\prime}}\geq C_{0}2^{k}$. By Theorem 5.3, we may choose $\beta$ sufficiently close to $\beta_{c}$ and $m,M,N$ large enough so that $\mathfrak{K}\neq\emptyset$. Introduce $\mathbf{U}:=\prod_{i=1}^{t}\mathbf{U}_{i}$, where
$\mathbf{U}_{i}:=\frac{1}{|\mathfrak{K}|}\sum_{k\in\mathfrak{K}}\frac{1}{A_{x_{i},y_{i}}(2^{k})}\sum_{u\in\mathbb{A}_{y_{i}}(2^{k})}\mathds{1}\{u\xleftrightarrow{\mathbf{n}_{i}+\mathbf{n}_{i}^{\prime}\>}x_{i}\},$ (5.4)
and,
$a_{x,y}(u):=\frac{\langle\sigma_{x}\sigma_{u}\rangle_{\beta}\langle\sigma_{u}\sigma_{y}\rangle_{\beta}}{\langle\sigma_{x}\sigma_{y}\rangle_{\beta}},\qquad A_{x,y}(p):=\sum_{u\in\mathbb{A}_{y}(p)}a_{x,y}(u).$ (5.5)
Using the switching lemma (2.4), we have that $\mathbf{E}_{\beta}^{\mathbf{xy},\emptyset}[\mathbf{U}]=1$. The following concentration inequality is a consequence of the definition of $\mathbb{A}_{y_{i}}(2^{k})$ and the properties of regular scales, and is still valid when $d=3$ (see [ADC21, Proposition 6.6]). The assumption on $m/n$ and $N/M$ below is justified by the definition of $\mathbb{A}_{y}(p)$.
Lemma 5.4 (Concentration of $\mathbf{U}$).
Assume that $x_{i}\in\Lambda_{n}$ and $y_{i}\notin\Lambda_{N}$ for all $1\leq i\leq t$. Additionally, assume that $m\geq 2dn$ and $N\geq 2dM$. Then, there exists $C=C(d,s)>0$ such that,
$\mathbf{E}_{\beta}^{\mathbf{xy},\emptyset}[(\mathbf{U}-1)^{2}]\leq\frac{C^{2}}{|\mathfrak{K}|}.$ (5.6)
We work under the assumptions of Lemma 5.4. Using the Cauchy–Schwarz inequality together with Lemma 5.4,
$\left|\mathbf{P}_{\beta}^{\mathbf{xy}}[E\cap F]-\mathbf{E}_{\beta}^{\mathbf{xy},\emptyset}\big{[}\mathbf{U}\cdot\mathds{1}\{(\mathbf{n}_{1},\ldots,\mathbf{n}_{s})\in E\cap F\}\big{]}\right|\leq\Big{(}\mathbf{E}_{\beta}^{\mathbf{xy},\emptyset}[(\mathbf{U}-1)^{2}]\Big{)}^{\tfrac{1}{2}}\leq\frac{C}{\sqrt{|\mathfrak{K}|}}.$ (5.7)
We now introduce the proper event which will allow us to “decouple” the events $E$ and $F$.
Definition 5.5.
Let $\mathbf{u}=(u_{1},\ldots,u_{t})$ with $u_{i}\in\textup{Ann}(m,M)$ for every $i$. The event $\mathcal{G}(u_{1},\ldots,u_{t})=\mathcal{G}(\mathbf{u})$ is defined as follows: for every $i\leq s$, there exists $\mathbf{k}_{i}\leq\mathbf{n}_{i}+\mathbf{n}_{i}^{\prime}$ such that $\mathbf{k}_{i}=0$ on $\Lambda_{n}$, $\mathbf{k}_{i}=\mathbf{n}_{i}+\mathbf{n}_{i}^{\prime}$ outside $\Lambda_{N}$, $\partial\mathbf{k}_{i}=\{u_{i},y_{i}\}$ for $i\leq t$, and $\partial\mathbf{k}_{i}=\emptyset$ for $t<i\leq s$.
By the switching principle (see [ADCTW19, Lemma 2.1]), one has
$\mathbf{P}_{\beta}^{\mathbf{x}\mathbf{y},\emptyset}\Big{[}\{(\mathbf{n}_{1},\ldots,\mathbf{n}_{s})\in E\cap F\}\cap\{u_{i}\xleftrightarrow{\mathbf{n}_{i}+\mathbf{n}_{i}^{\prime}\>}y_{i},\>\forall 1\leq i\leq t\}\cap\>\mathcal{G}(\mathbf{u})\Big{]}\\ =\Big{(}\prod_{i=1}^{t}a_{x_{i},y_{i}}(u_{i})\Big{)}\mathbf{P}_{\beta}^{\mathbf{xu},\mathbf{uy}}\left[(\mathbf{n}_{1},\ldots,\mathbf{n}_{s})\in E,(\mathbf{n}_{1}^{\prime},\ldots,\mathbf{n}_{s}^{\prime})\in F,\>\mathcal{G}(\mathbf{u})\right].$ (5.8)
The identity,
$\mathbf{P}_{\beta}^{\mathbf{xu},\mathbf{uy}}[(\mathbf{n}_{1},\ldots,\mathbf{n}_{s})\in E,(\mathbf{n}_{1}^{\prime},\ldots,\mathbf{n}_{s}^{\prime})\in F]\\ =\mathbf{P}_{\beta}^{\mathbf{xu}}[(\mathbf{n}_{1},\ldots,\mathbf{n}_{s})\in E]\mathbf{P}_{\beta}^{\mathbf{uy}}[(\mathbf{n}_{1}^{\prime},\ldots,\mathbf{n}_{s}^{\prime})\in F]$ (5.9)
motivates us to prove that under $\mathbf{P}_{\beta}^{\mathbf{xu},\mathbf{uy}}$, the event $\mathcal{G}(\mathbf{u})$ occurs with high probability. This is the technical step in the proof of [ADC21] that requires some work to be extended to the three-dimensional case. Recall that $n\leq m\leq M\leq N$.
Lemma 5.6.
We keep the assumptions of Lemma 5.4. Assume $\tfrac{\beta_{c}}{2}\leq\beta\leq\beta_{c}$. For any fixed value $\alpha=m/M$, there exist $m,N>0$ large enough (and which only depend on $\varepsilon,\alpha,n,s,d$) such that, for every $\mathbf{u}$ with $u_{i}\in\mathbb{A}_{y_{i}}(2^{k_{i}})$ with $k_{i}\in\mathfrak{K}$ for $1\leq i\leq t$,
$\mathbf{P}_{\beta}^{\mathbf{x}\mathbf{y},\emptyset}\Big{[}\{u_{i}\xleftrightarrow{\mathbf{n}_{i}+\mathbf{n}_{i}^{\prime}\>}y_{i},\>\forall 1\leq i\leq t\}\cap\>\mathcal{G}(\mathbf{u})\Big{]}\Big{(}\prod_{i=1}^{t}a_{x_{i},y_{i}}(u_{i})\Big{)}^{-1}=\mathbf{P}_{\beta}^{\mathbf{xu},\mathbf{uy}}[\mathcal{G}(\mathbf{u})^{c}]\leq\frac{\varepsilon}{2}.$ (5.10)
Proof.
Below, the constants $C_{i}$ only depend on $d$ and $s$. The equality follows from an application of the switching lemma, we thus focus on the inequality. Write $\mathcal{G}(\mathbf{u})=\cap_{1\leq i\leq s}G_{i}$ (where the definition of $G_{i}$ is implicit). Then, $H_{i}^{c}\cap F_{i}^{c}\subset G_{i}$ where,
$H_{i}:=\{\text{Ann}(M,N)\text{ is crossed by a cluster in }\mathbf{n}_{i}\},$ (5.11)
and
$F_{i}:=\{\text{Ann}(n,m)\text{ is crossed by a cluster in }\mathbf{n}_{i}^{\prime}\}.$ (5.12)
Indeed, if $H_{i}^{c}\cap F_{i}^{c}$ occurs, we may define $\mathbf{k}_{i}$ as the sum of the restriction of $\mathbf{n}_{i}$ to the clusters intersecting $\Lambda_{N}^{c}$ and the restriction of $\mathbf{n}_{i}^{\prime}$ to the clusters intersecting $\Lambda_{m}^{c}$. Introduce an intermediate scale $n\leq r\leq m$. Using a union bound,
$\mathbf{P}_{\beta}^{\mathbf{xu},\mathbf{uy}}[\mathcal{G}(\mathbf{u})^{c}]\leq\sum_{i=1}^{t}\Big{(}\mathbf{P}_{\beta}^{x_{i}u_{i}}[H_{i}]+\mathbf{P}_{\beta}^{u_{i}y_{i}}[F_{i}]\Big{)}+(s-t)\Big{(}\mathbf{P}_{\beta}^{\emptyset}[H_{s}]+\mathbf{P}_{\beta}^{\emptyset}[F_{s}]\Big{)}.$ (5.13)
We progressively fix $r,m,N$ large enough (recall that $m/M=\alpha$ is fixed).
Bound on $\mathbf{P}_{\beta}^{u_{i}y_{i}}[F_{i}]$
Fix $1\leq i\leq t$. Notice that
$\mathbf{P}_{\beta}^{u_{i}y_{i}}[F_{i}]\leq\mathbf{P}_{\beta}^{u_{i}y_{i}}[\Gamma(\mathbf{n}_{i}^{\prime})\textup{ crosses }\textup{Ann}(r,m)]+\mathbf{P}_{\beta}^{u_{i}y_{i}}[\mathbf{n}_{i}^{\prime}\setminus\overline{\Gamma}(\mathbf{n}_{i}^{\prime})\textup{ crosses }\textup{Ann}(n,r)],$ (5.14)
see Figure 4. Now, using the chain rule (Proposition 4.6),
$\mathbf{P}_{\beta}^{u_{i}y_{i}}[\Gamma(\mathbf{n}_{i}^{\prime})\textup{ crosses }\textup{Ann}(r,m)]$ $\leq\sum_{v\in\partial\Lambda_{r}}\frac{\langle\sigma_{u_{i}}\sigma_{v}\rangle_{\beta}\langle\sigma_{v}\sigma_{y_{i}}\rangle_{\beta}}{\langle\sigma_{u_{i}}\sigma_{y_{i}}\rangle_{\beta}}$ (5.15)
$\leq C_{1}\frac{r^{2}}{m}\max_{v\in\partial\Lambda_{r}}\frac{\langle\sigma_{v}\sigma_{y_{i}}\rangle_{\beta}}{\langle\sigma_{u_{i}}\sigma_{y_{i}}\rangle_{\beta}}$ (5.16)
$\leq C_{2}\frac{r^{2}}{m},$ (5.17)
where we used the fact that $u\in\mathbb{A}_{y_{i}}(2^{k_{i}})$ on the third line.
Figure 4: An illustration of a configuration $\mathbf{n}_{i}^{\prime}$ contributing to the event $F_{i}$ introduced in (5.12). The backbone of $\mathbf{n}_{i}^{\prime}$ is the bold black line. Here, $\Gamma(\mathbf{n}_{i}^{\prime})$ does not cross $\textup{Ann}(r,m)$ but $\mathbf{n}_{i}^{\prime}$ crosses $\textup{Ann}(n,m)$ thanks to a string of loops crossing $\textup{Ann}(n,r)$.
Then, arguing as in (4.50) (see also the proof of[Pan23, Lemma 6.3]),
$\mathbf{P}_{\beta}^{u_{i}y_{i}}[\mathbf{n}_{i}^{\prime}\setminus\overline{\Gamma}(\mathbf{n}_{i}^{\prime})\textup{ crosses }\textup{Ann}(n,r)]=\sum_{\gamma:u_{i}\rightarrow y_{i}}\mathbf{P}_{\beta}^{u_{i}y_{i}}[\Gamma(\mathbf{n}_{i}^{\prime})=\gamma]\mathbf{P}^{\emptyset}_{\overline{\gamma}^{c},\beta}[\partial\Lambda_{n}\xleftrightarrow{\mathbf{n}_{i}^{\prime}\>}\partial\Lambda_{r}].$ (5.18)
However, by Proposition 2.5, one has
$\mathbf{P}^{\emptyset}_{\overline{\gamma}^{c},\beta}[\partial\Lambda_{n}\xleftrightarrow{\mathbf{n}_{i}^{\prime}\>}\partial\Lambda_{r}]$ $\leq\phi_{\overline{\gamma}^{c},\beta_{c}}^{0}[\partial\Lambda_{n}\xleftrightarrow{}\partial\Lambda_{r}]$ (5.19)
$\leq\phi_{\beta_{c}}[\partial\Lambda_{n}\xleftrightarrow{}\partial\Lambda_{r}],$ (5.20)
where we additionally used monotonicity properties of the FK-Ising measure (see [Gri06, DC19]). Using Theorem 5.1, we find that for $r$ large enough, one has
$\phi_{\beta_{c}}[\partial\Lambda_{n}\xleftrightarrow{}\partial\Lambda_{r}]\leq\frac{\varepsilon}{8s}.$ (5.21)
Combining this observation and (5.17), we obtain that for $m$ large enough,
$\mathbf{P}_{\beta}^{u_{i}y_{i}}[F_{i}]\leq\frac{\varepsilon}{4s}.$ (5.22)
Bound on $\mathbf{P}_{\beta}^{\emptyset}[H_{s}]$
Using Proposition 2.5,
$\mathbf{P}_{\beta}^{\emptyset}[H_{s}]\leq\phi_{\beta_{c}}[\partial\Lambda_{n}\xleftrightarrow{}\partial\Lambda_{m}].$ (5.23)
Using Theorem 5.1 and choosing $m$ large enough (and such that (5.22) holds) therefore yields
$\mathbb{P}_{\beta}^{\emptyset}[H_{s}]\leq\frac{\varepsilon}{4s}$ (5.24)
Bound on $\mathbf{P}_{\beta}^{x_{i}u_{i}}[H_{i}]$
Fix $1\leq i\leq t$. Fix $r,m$ such that the above bounds holds. Notice that this fixes the value of $M$ by assumption. Still using Proposition 2.5, one has
$\mathbf{P}_{\beta}^{x_{i}u_{i}}[H_{i}]=\phi_{\beta}[H_{i}\>|\>x_{i}\xleftrightarrow{}u_{i}]\leq\max_{w,v\in\Lambda_{m}}\phi_{\beta_{c}/2}[w\xleftrightarrow{}v]^{-1}\cdot\phi_{\beta_{c}}[\partial\Lambda_{M}\xleftrightarrow{}\partial\Lambda_{N}],$ (5.25)
which can be made smaller than $\tfrac{\varepsilon}{4s}$ by using Theorem 5.1 and choosing $N>0$ large enough.
Bound on $\mathbf{P}_{\beta}^{\emptyset}[F_{s}]$
We notice that
$\mathbf{P}_{\beta}^{\emptyset}[F_{s}]\leq\phi_{\beta_{c}}[\partial\Lambda_{M}\xleftrightarrow{}\partial\Lambda_{N}],$ (5.26)
which we bound similarly by $\tfrac{\varepsilon}{4s}$ for $N$ large enough.
Plugging the above bounds in (5.13), we obtain
$\mathbf{P}_{\beta}^{\mathbf{xu},\mathbf{uy}}[\mathcal{G}(\mathbf{u})^{c}]\leq\sum_{i=1}^{t}\Big{(}\frac{\varepsilon}{4s}+\frac{\varepsilon}{4s}\Big{)}+(s-t)\Big{(}\frac{\varepsilon}{4s}+\frac{\varepsilon}{4s}\Big{)}\leq\frac{\varepsilon}{2},$ (5.27)
which concludes the proof. ∎
We can now conclude.
Proof of Theorem 2.4.
Introduce the weights $\delta(\mathbf{u},\mathbf{x},\mathbf{y})$ defined by
$\delta(\mathbf{u},\mathbf{x},\mathbf{y}):=\mathds{1}\{\exists(k_{1},\ldots,k_{t})\in\mathfrak{K}^{t},\>\mathbf{u}\in\mathbb{A}_{y_{1}}(2^{k_{1}})\times\ldots\times\mathbb{A}_{y_{t}}(2^{k_{t}})\}\prod_{i=1}^{t}\frac{a_{x_{i},y_{i}}(u_{i})}{|\mathfrak{K}|A_{x_{i},y_{i}}(2^{k_{i}})}.$ (5.28)
Notice that
$\sum_{\begin{subarray}{c}(k_{1},\ldots,k_{t})\in\mathfrak{K}^{t}\\ \mathbf{u}\in\mathbb{A}_{y_{1}}(2^{k_{1}})\times\ldots\times\mathbb{A}_{y_{t}}(2^{k_{t}})\end{subarray}}\delta(\mathbf{u},\mathbf{x},\mathbf{y})=1.$ (5.29)
We abbreviate the above sum as $\sum_{\mathbf{u}}$. Assume that $m/M=\alpha$. Let $\alpha>0$ to be chosen small enough, and let $m=m(\alpha),N=N(\alpha)$ be given by Lemma 5.6.
Let $\mathbf{x}=(x_{1},\ldots,x_{t})\in(\Lambda_{n})^{t}$ and $\mathbf{y}=(y_{1},\ldots,y_{t})\in(\Lambda_{N}^{c})^{t}$. Equation (5.7) together with Lemma 5.6 yield
$\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E\cap F]-\sum_{\mathbf{u}}\delta(\mathbf{u},\mathbf{x},\mathbf{y})\mathbf{P}_{\beta}^{\mathbf{xu}}[E]\mathbf{P}_{\beta}^{\mathbf{uy}}[F]\Big{|}\\ \leq\frac{C}{\sqrt{|\mathcal{K}|}}+\frac{\varepsilon}{2}.$ (5.30)
We now use Theorem 5.3 to argue that if $\beta$ is sufficiently close to $\beta_{c}$ and if $M/m$ is large enough (which affects how large $N$ is), then $|\mathfrak{K}|\geq 2\varepsilon^{-2}C$. This gives for every $\mathbf{y}\in(\Lambda_{N}^{c})^{t}$,
$\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E\cap F]-\sum_{\mathbf{u}}\delta(\mathbf{u},\mathbf{x},\mathbf{y})\mathbf{P}_{\beta}^{\mathbf{xu}}[E]\mathbf{P}_{\beta}^{\mathbf{uy}}[F]\Big{|}\\ \leq\varepsilon.$ (5.31)
Proof of (2.9)
We begin by proving (2.9) when $y_{i},y_{i}^{\prime}$ are in regular scales (but not necessarily the same ones). Applying the above inequality once for $\mathbf{y}$ and once for $\mathbf{y^{\prime}}$ with the event $E$ and $F=\Omega_{\mathbb{Z}^{d}}$,
$\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E]-\mathbf{P}_{\beta}^{\mathbf{xy^{\prime}}}[E]\Big{|}\leq\Big{|}\sum_{\mathbf{u}}(\delta(\mathbf{u},\mathbf{x},\mathbf{y})-\delta(\mathbf{u},\mathbf{x},\mathbf{y^{\prime}}))\mathbf{P}_{\beta}^{\mathbf{xu}}[E]\Big{|}+2\varepsilon.$ (5.32)
Since all the $y_{i},y_{i}^{\prime}$ are in regular scales, one has $\mathbb{A}_{y_{i}}(2^{k_{i}})=\mathbb{A}_{y_{i}^{\prime}}(2^{k_{i}})=\text{Ann}(2^{k_{i}},2^{k_{i}+1}).$ Moreover, using Property $\mathbf{(P2)}$ of regular scales, there exists $C_{1}=C_{1}(s,d)>0$ such that
$\left|\delta(\mathbf{u},\mathbf{x},\mathbf{y})-\delta(\mathbf{u},\mathbf{x},\mathbf{y^{\prime}})\right|\leq C_{1}\Big{(}\frac{M}{N}\Big{)}\delta(\mathbf{u},\mathbf{x},\mathbf{y})\leq\varepsilon\delta(\mathbf{u},\mathbf{x},\mathbf{y}),$ (5.33)
if $N$ is large enough. Indeed, in this configuration $\delta(\mathbf{u},\mathbf{x},\mathbf{y})$ and $\delta(\mathbf{u},\mathbf{x},\mathbf{y}^{\prime})$ are both close to
$\prod_{i\leq t}\frac{\langle\sigma_{x_{i}}\sigma_{u_{i}}\rangle}{|\mathfrak{K}|\sum_{v_{i}\in\textup{Ann}(2^{k_{i}},2^{k_{i}+1})}\langle\sigma_{x_{i}}\sigma_{v_{i}}\rangle}.$ (5.34)
This gives that for every $\mathbf{y},\mathbf{y}^{\prime}\in(\Lambda_{N}^{c})^{t}$ having all coordinates in regular scales,
$\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E]-\mathbf{P}_{\beta}^{\mathbf{xy^{\prime}}}[E]\Big{|}\leq 3\varepsilon.$ (5.35)
Note that we can increase $m,M,N$, while maintaining the value $M/m$, and choose $\beta$ even closer to $\beta_{c}$, to also obtain (using the exact same argument): for every $\mathbf{z},\mathbf{z}^{\prime}\in(\Lambda_{m}^{c})^{t}$ having all coordinates in regular scales,
$\Big{|}\mathbf{P}_{\beta}^{\mathbf{xz}}[E]-\mathbf{P}_{\beta}^{\mathbf{xz^{\prime}}}[E]\Big{|}\leq 3\varepsilon.$ (5.36)
We now consider $\mathbf{z}=(z_{1},\ldots,z_{t})$ with $z_{i}\in\text{Ann}(m,M)$ in a regular scale. Also, pick $\mathbf{y}$ on which we do not assume anything. We have,
$\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E]-\mathbf{P}_{\beta}^{\mathbf{xz}}[E]\Big{|}$ $=$ $\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E]-\sum_{\mathbf{u}}\delta(\mathbf{u},\mathbf{x},\mathbf{y})\mathbf{P}_{\beta}^{\mathbf{xz}}[E]\Big{|}$
$\leq$ $\Big{|}\mathbf{P}_{\beta}^{\mathbf{xy}}[E]-\sum_{\mathbf{u}}\delta(\mathbf{u},\mathbf{x},\mathbf{y})\mathbf{P}_{\beta}^{\mathbf{xu}}[E]\Big{|}+3\varepsilon$
$\leq$ $4\varepsilon,$
where on the second line we used (5.36) with $(\mathbf{y},\mathbf{y}^{\prime})=(\mathbf{u},\mathbf{z})$, and in the third line we used (5.31) with $F=\Omega_{\mathbb{Z}^{d}}$. This gives (2.9).
Proof of (2.10)
The same argument works for (2.10) for every $\mathbf{x},\mathbf{x^{\prime}},\mathbf{y}$, noticing that for every regular $\mathbf{u}$ for which $\delta(\mathbf{u},\mathbf{x},\mathbf{y})\neq 0$, using once again $\mathbf{(P2)}$,
$\left|\delta(\mathbf{u},\mathbf{x},\mathbf{y})-\delta(\mathbf{u},\mathbf{x^{\prime}},\mathbf{y})\right|\leq C_{2}\left(\frac{n}{m}\right)\delta(\mathbf{u},\mathbf{x},\mathbf{y})\leq\varepsilon\delta(\mathbf{u},\mathbf{x},\mathbf{y})$ (5.37)
if we take $m$ sufficiently large.
Proof of (2.8)
To obtain (2.8) we repeat the same line of reasoning. We start by applying (5.31). Then, we replace each $\mathbf{P}_{\beta}^{\mathbf{xu}}[E]$ by $\mathbf{P}_{\beta}^{\mathbf{xy}}[E]$ using (5.36). Similarly, we replace $\mathbf{P}_{\beta}^{\mathbf{uy}}[F]$ by $\mathbf{P}_{\beta}^{\mathbf{xy}}[F]$. The proof follows readily. ∎
6 Proof of Proposition 3.2
In this last section, we use the techniques introduced in Section 5 to prove Proposition 3.2.
Proof of Proposition 3.2.
We begin with $(i)$. The argument follows from the fact that the IIC of the measure of interest can be described by the union of the backbone and a collection of finite loops. If the IIC has two ends, then there must be an infinite path which does not use any edges of the backbone. Fix $y\in\mathbb{Z}^{d}$, $n\geq 1$ and $|x|\gg 1$. Reasoning again as in (4.49),
$\mathbf{P}^{0x,\emptyset}_{\beta_{c}}[y\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\>}\partial\Lambda_{n}]\leq\sum_{u\in\partial\Lambda_{n}}\langle\sigma_{u}\sigma_{y}\rangle_{\beta_{c}}^{2}\leq\frac{C_{3}}{n^{d-3}}.$ (6.1)
Hence, by Theorem 3.1,
$\mathbf{P}^{0\infty,\emptyset}[y\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\>}\partial\Lambda_{n}]\leq\frac{C_{3}}{n^{d-3}},$ (6.2)
which immediately yields,
$\mathbf{P}^{0\infty,\emptyset}[y\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\>}\infty]=0$ (6.3)
and the result.
We turn to $(ii)$. Notice that by the switching lemma, if $x,y\in\mathbb{Z}^{d}$,
$\mathbf{P}_{\beta_{c}}^{0x,\emptyset}[0\xleftrightarrow{}y]=\frac{\langle\sigma_{0}\sigma_{y}\rangle_{\beta_{c}}\langle\sigma_{y}\sigma_{x}\rangle_{\beta_{c}}}{\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}}.$ (6.4)
The gradient estimate [ADC21, Proposition 5.9] gives that
$\lim_{|x|\rightarrow\infty}\frac{\langle\sigma_{y}\sigma_{x}\rangle_{\beta_{c}}}{\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}}=1,$ (6.5)
which seems to imply (3.10) by Theorem 3.1. However, the event $\{0\xleftrightarrow{}y\}$ is not local [^14] For each fixed $x$, this event can be approximated by local events using the fact that the measure $\mathbf{P}_{\beta_{c}}^{0x,\emptyset}$ does not percolate, see[ADCS15]. Here we need a version of that statement which is uniform in $x$. This is why the qualitative result of [ADCS15] is not enough when $d=3$., so we need to work a little to conclude. Let $2|y|\leq m\leq M$ to be taken large enough. We follow the argument used in the proof of Lemma 5.6 and notice that
$\mathbf{P}_{\beta_{c}}^{0x,\emptyset}[\{0\xleftrightarrow{\Lambda_{M}}y\}^{c}\cap\{0\xleftrightarrow{}y\}]\leq\mathbf{P}_{\beta_{c}}^{0x}[\mathcal{H}_{1}]+\mathbf{P}_{\beta_{c}}^{0x,\emptyset}[\mathcal{H}_{2}],$ (6.6)
where $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are two events defined as follows:
  1. -
    $\mathcal{H}_{1}$: the backbone $\Gamma(\mathbf{n}_{1})$ does two successive crossings of $\textup{Ann}(m,M)$,
  2. -
    $\mathcal{H}_{2}$: the event $\mathcal{H}_{1}$ does not occur and there exists $s\in\Lambda_{m}$ such that $y\xleftrightarrow{\mathbf{n}_{1}+\mathbf{n}_{2}\setminus\overline{\Gamma}(\mathbf{n}_{1})\>}s$.
Using the chain rule for backbones and (1.11),
$\mathbf{P}_{\beta_{c}}^{0x}[\mathcal{H}_{1}]\leq\sum_{\begin{subarray}{c}u\in\partial\Lambda_{m}\\ v\in\partial\Lambda_{M}\end{subarray}}\frac{\langle\sigma_{0}\sigma_{v}\rangle_{\beta_{c}}\langle\sigma_{v}\sigma_{u}\rangle_{\beta_{c}}\langle\sigma_{u}\sigma_{x}\rangle_{\beta_{c}}}{\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}}\leq\max_{u\in\Lambda_{m}}\frac{\langle\sigma_{u}\sigma_{x}\rangle_{\beta_{c}}}{\langle\sigma_{0}\sigma_{x}\rangle_{\beta_{c}}}\cdot C_{1}\frac{m^{d-1}}{M^{d-3}},$ (6.7)
where $C_{1}=C_{1}(d)>0$. Moroever, a similar reasoning a in (4.49) yields the existence of $C_{2}=C_{2}(d)>0$ such that,
$\mathbf{P}_{\beta_{c}}^{0x,\emptyset}[\mathcal{H}_{2}]\leq\sum_{u\in\partial\Lambda_{m}}\langle\sigma_{u}\sigma_{y}\rangle_{\beta_{c}}^{2}\leq\frac{C_{2}}{m^{d-3}},$ (6.8)
where we used (1.11) in the last inequality. Taking $m=M^{1/4}$ and using the gradient estimate [ADC21, Proposition 5.9] again,
$\limsup_{M\rightarrow\infty}\limsup_{|x|\rightarrow\infty}\mathbf{P}_{\beta_{c}}^{0x,\emptyset}[\{0\xleftrightarrow{\Lambda_{M}}y\}^{c}]=0.$ (6.9)
The proof of (3.10) follows readily. Moreover, (3.11) follows from (1.11) and [DCP24, Theorem 1.4].
Acknowledgements.
We thank Hugo Duminil-Copin for inspiring discussions and for constant support. We thank Gordon slade for useful comments leading to the discussion below Theorem 1.4 and to Proposition 3.2. We also thank Lucas D’Alimonte, Trishen S. Gunaratnam, Yucheng Liu, Mathilde Pacholski and Alexis Prévost for useful comments. This project has received funding from the Swiss National Science Foundation, the NCCR SwissMAP, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 757296).